### v1

# POLAR: A Benchmark for Multilingual, Multicultural, and Multi-Event Online Polarization

Abstract
Online polarization poses a growing challenge for democratic discourse, yet most computational social science research remains monolingual, culturally narrow, or event-specific. We introduce POLAR, a multilingual, multicultural, and multievent dataset with over 23k instances in seven languages from diverse online platforms and real-world events. Polarization is annotated along three axes: presence, type, and manifestation, using a variety of annotation platforms adapted to each cultural context. We conduct two main experiments: (1) we fine-tune six multilingual pretrained language models in both monolingual and cross-lingual setups; and (2) we evaluate a range of open and closed large language models (LLMs) in few-shot and zero-shot scenarios. Results show that while most models perform well on binary polarization detection, they achieve substantially lower scores when predicting polarization types and manifestations. These findings highlight the complex, highly contextual nature of polarization and the need for robust, adaptable approaches in NLP and computational social science. All resources will be released to support further research and effective mitigation of digital polarization globally.
POLAR: A Benchmark for Multilingual, Multicultural, and Multi-Event Online Polarization
Usman Naseem1, Juan Ren1, Saba Anwar2, Sarah Kohail6, Rudy Alexandro Garrido Veliz2, Robert Geislinger2, Aisha Jabr6, Idris Abdulmumin5, Laiba Qureshi2, Aarushi Ajay Borkar2, Maryam Ibrahim Mukhtar7, Abinew Ali Ayele2,3, Ibrahim Said Ahmad7,8, Adem Ali2,3, Martin Semmann2, Shamsuddeen Hassan Muhammad4,7, Seid Muhie Yimam2 1Macquarie University, 2University of Hamburg, 3Bahir Dar University,4Imperial College London 5University of Pretoria, 6Zayed University, 7Bayero University Kano, 8Northeastern University
Figure 1: Pipeline for POLAR construction: multi-platform data curation in 7 languages, annotation workflow with quality control, and benchmarking.
1 Introduction
Online polarization, defined as sharp division and antagonism between social, political, or identity groups, has become a pervasive threat to democratic institutions, civil discourse, and social cohesion worldwide (Waller and Anderson, 2021; Iandoli et al., 2021). It is often fueled by biased or inflammatory content on social media, reinforcing echo chambers and undermining mutual understanding (Garimella, 2018). Polarized discourse not only amplifies ideological divides but can also escalate into hate speech, harassment, and real-world violence. As such, early detection of polarization is critical to designing interventions that promote healthier online ecosystems.
Despite growing attention, computational approaches to polarization suffer from major limitations. First, most existing datasets focus on English or high-resource languages, reflecting a widespread trend across NLP tasks that ignores the rich diversity of linguistic and sociocultural contexts in which polarization manifests. Second, current benchmarks are often event-specific or monodomain, such as U.S. elections or Western political debates, limiting their generalizability. Third, the conceptualization of polarization in NLP has largely been binary or topic-focused, overlooking the multifaceted ways in which polarization is expressed through vilification, dehumanization, stereotyping, or other rhetorical tactics.
To address these gaps, we introduce POLAR a novel multilingual, multicultural, and multievent dataset for fine-grained polarization detection. It spans seven languages across diverse regions, including low-resource languages such as Amharic and Hausa. Our data is sourced from various platforms (e.g., Twitter/X, Facebook, BlueSky, Reddit, and local news outlets), reflecting authentic, event-driven discourse ranging from armed conflict (e.g., the Tigray War) to social justice movements (e.g., abortion rights, migration crises).
Unlike prior work, POLAR supports three complementary tasks:
  1. 1.
    Binary Polarization Detection: Is a message polarized or not?
  2. 2.
    Polarization Type Classification: What social dimension is targeted (e.g., political, religious, racial)?
  3. 3.
    Manifestation Identification: How is polarization rhetorically expressed (e.g., stereotyping, deindividuation, extreme language)?
We design a cross-cultural annotation protocol tailored for each language’s sociopolitical context. Extensive qualitative analysis reveals how event salience, linguistic norms, and platform affordances shape polarization dynamics across languages. The complete pipeline can be found in Figure 1. We benchmark a range of multilingual language models (MLMs) and large language models (LLMs) under zero-shot, few-shot, and cross-lingual scenarios. Our experiments highlight the challenges of generalization and the limitations of current models in capturing nuanced rhetorical patterns across languages. Our contributions are as follows:
  • We release POLAR, the first large-scale, multilingual, fine-grained dataset for polarization detection across 7 languages and diverse global events.
  • We define a taxonomy of polarization types and manifestations, operationalized through a robust cross-lingual annotation protocol.
  • We provide comprehensive benchmarks using state-of-the-art MLMs and LLMs across multiple evaluation settings (monolingual, cross-lingual, and few-shot).
Figure 2: Languages represented in the dataset, covering diverse linguistic and regional contexts.
2 Related Work
Online polarization has long been recognized as a threat to democracy and social cohesion, intensifying through social media echo chambers and biased content (Waller and Anderson, 2021; Iandoli et al., 2021; Garimella, 2018). As social media and other online platforms become key arenas for political and cultural discourse, the need for early detection and nuanced understanding of polarization has grown significantly. Such efforts are critical not only for content moderation, but also for peacebuilding, policy development, and responsible digital governance. Foundational research has defined polarization as both intergroup hostility and ingroup cohesion (Arora et al., 2022), and has highlighted its relationship with hate speech, fragmentation, and incivility (Bozdag and van den Hoven, 2020; Mathew et al., 2021).
A growing body of research has documented the role of online spaces in intensifying polarization across different regions (Kubin and von Sikorski, 2021; Barberá, 2020; Gitlin, 2016; Soares and Recuero, 2021). However, most computational work focuses on high-resource languages and event- or region-specific datasets, limiting generalizability (Kubin and von Sikorski, 2021). This leaves a significant gap in our ability to generalize findings across cultures, languages, and events, especially in the Global South or multilingual regions.
The lack of standardized datasets across languages has hindered progress in developing and evaluating polarization detection models with cross-lingual or cross-cultural capabilities. Recent shared tasks on hate speech and toxicity (Basile and others, 2019; Mohammad and others, 2021; Pamungkas et al., 2020; Ousidhoum et al., 2024) have expanded the language and domain coverage, yet remain less fine-grained regarding polarization’s diverse types and rhetorical manifestations. Our work addresses this gap by presenting the first comprehensive, fine-grained dataset benchmark for multilingual, multicultural, and multievent online polarization, enabling robust cross-lingual and context-aware modeling.
3 POLAR Dataset Construction
3.1 Data Collection
We collected data from various online platforms, including X (formerly Twitter), Facebook, Reddit, Bluesky, Threads, and news/commentary forums. We use a dynamic keyword-driven strategy tailored for each language. Human experts curated keyword lists to reflect culturally and politically significant discourse across regions and events. Table 1 shows the languages covered, data splits and total number of instances annotated for each language.
The Amharic dataset focuses on the Northern Ethiopia or Tigray War in Ethiopia; Arabic texts cover a broad array of social and economic topics; English data centers on US elections and international conflicts; the German corpus features election discourse and migration debates; the Hausa subset captures religious and ideological discussions in West and Central Africa; Spanish content spans abortion, migration, and indigenous/gender rights; and Urdu texts reflect political and sectarian divides.
Language Source(s) Train Dev Test Total
Amharic Facebook, X 3500 500 1000 5000
Arabic Facebook, X, Threads, News 1482 212 424 2118
English X, BlueSky, News 2117 303 605 3025
German X, BlueSky, Reddit 2426 347 694 3467
Hausa Facebook , X 3893 557 1113 5563
Spanish X, BlueSky 1400 201 401 2002
Urdu X 1960 280 560 2800
Table 1: Dataset sources and split sizes for all three tasks as per each language.
3.2 Annotation Process
Given the cultural and linguistic breadth of POLAR , we developed detailed, multilingual annotation guidelines and deployed a hybrid strategy combining crowdsourcing with trained community annotators.
Annotation Guidelines: We developed the guidelines in English and Amharic, and then translated and culturally adapted them for each target language. Annotators were instructed to:
  • Identify whether a text is polarized;
  • If the text is classified as polarized, tag the type of polarization (political, racial/ethnic, religious, gender/sexual identity, other);
  • If the text is classified as polarized, tag its manifestations (stereotyping, vilification, dehumanization, deindividuation, extreme language, lack of empathy, invalidation).
Multiple labels were allowed due to the conceptual and contextual overlap often observed in polarized content.
Annotator Recruitment and Annotation Process
We used both expert and crowd-sourced annotation strategies. Trained annotators used POTATO and Label Studio, while crowdsourced annotators used Mechanical Turk and Prolific. To evaluate inter-annotator agreement, we report both Cohen’s Kappa and Fleiss’ Kappa. Since different teams were responsible for dataset creation across various languages, we provide a detailed description of the annotation process for each language:
Language Total Polarization Polarization Types Polarization Manifestations
gender/sexual political religious racial/ethnic other vilification extreme_language stereotype invalidation lack_of_empathy dehumanization
Amharic 5000 3753 29 3342 99 1297 189 2398 1527 2729 799 880 657
Arabic 2118 338 102 178 86 171 220 413 341 283 240 175 102
English 3025 1021 18 892 44 98 5 342 179 138 85 45 39
German 3467 1596 241 1981 448 782 653 1371 1001 1420 1258 1124 568
Hausa 5563 625 46 285 149 183 22 74 179 253 13 53 202
Spanish 2002 1057 254 742 571 521 405 798 622 843 205 599 307
Urdu 2800 1860 236 1533 557 423 126 1548 1409 979 670 632 501
Table 2: Number of samples labeled positive for each annotation task across languages. Labels are grouped by task: Polarization (Task 1), Polarization Types (Task 2), and Polarization Manifestations (Task 3).
  • Amharic: We used train annotators to annotated 5,000 samples in five batches using the POTATO annotation tool. Annotators completed three rounds of guideline training and received continuous feedback, with ongoing supervision to ensure quality control. The annotation quality achieved a Fleiss’ Kappa of 0.49, with full agreement in 69.62% of cases and a peak pairwise Cohen’s Kappa of 0.65.
  • Arabic: Three topic-aware annotators labeled 2,118 samples using POTATO. Final polarization distribution was 16% positive, with Cohen’s Kappa at 0.296. Annotations were revised to ensure thematic consistency.
  • German: Annotation was conducted through Prolific. Annotators were screened via surveys and underwent pilot testing. Final annotation used a locally hosted POTATO interface adapted for multilabel inputs. Inter-annotator agreement was moderate, with notable consistency on polarized cases.
  • Hausa: We used trained annotators and Label Studio with iterative feedback and internal consistency checks. Data focused on religion and ideology on social media.
  • Spanish and English: Amazon Mechanical Turk (MTurk) was used with strict quality filters, such as task completion history and approval ratings. Annotators received feedback and were contacted throughout for quality assurance. English annotation reached a Kappa of 0.52 (up from 0.31 after worker filtering); Spanish showed lower agreement with an average Kappa of 0.24.
  • Urdu: Two approaches were used: (1) Manual Annotation: 1,792 samples labeled in batches with average Cohen’s Kappa of 0.55 (batch 1: 0.79; batch 2: 0.56; batch 3: 0.39). Disagreements were resolved by a fifth annotator. (2) Prolific: 1,460 samples were annotated in three batches. Initial agreement was low (Fleiss’ Kappa = 0.05). Additional filters (education, region) raised agreement in later batches to 0.28 and 0.30. The final dataset contains 2,800 high-quality Urdu samples.
Figure 3: Percentage of polarization per language
(a) Types
(b) Manifestation
Figure 4: Annotation statistics for polarization classification, types, and manifestations
3.3 Dataset Characteristics
Polarization Prevalence: As shown in Figure 3, Amharic (75%) and Urdu (66%) has the highest polarization rates, tied to their sources - Northern Ethiopia’s Tigray conflict and Pakistan’s political and sectarian climate. Spanish (53%), German (46%), and English (34%) fall in the mid-range. Arabic (12%) and Hausa (11%) reflect broader topical focus and possibly platform moderation effects. Hausa shows the lowest polarization rate despite covering sensitive religious and ideological topics, possibly due to moderation or cultural norms in expression. Datasets with acute sociopolitical events (e.g., Amharic, Urdu) show higher polarization. Topic scope, moderation, and platform dynamics significantly influence polarization prevalence.
Types of Polarization: Figure 4 show types of polarization. Political polarization dominates in Amharic (67%), German (57%), Urdu (55%), and Spanish (37%). Racial/ethnic and religious types also appear prominently in Spanish, Amharic, and German. Gender/sexual identity polarization is generally low but visible in Spanish (13%) and Urdu (8%). Thematic emphasis (e.g., migration, elections) shapes type distribution. Spanish and Urdu show diverse forms of identity-based polarization, while political conflict dominates Amharic, Urdu, and German.
Polarization Manifestations: Figure 4 show types of Manifestation. Stereotyping and vilification are most frequent in Urdu, Amharic, German, and Spanish. Urdu’s vilification rate is highest (55%). Extreme language is strong in Urdu (50%), Spanish (31%), and Amharic (31%). Invalidation and dehumanization peak in German and Urdu. Conflict-heavy datasets (Urdu, Amharic) correlate with higher rates of hostile manifestations. Platform moderation and topic scope likely suppress such features in Arabic and Hausa.
Together, these analyses highlight how linguistic, cultural, and contextual variables - alongside event salience and platform dynamics - shape the structure and tone of online polarization. The POLAR dataset (see Table 2 for detailed statistics of POLAR dataset) offers a robust benchmark for examining these effects at scale.
4 Experimentation and Results
4.1 Experimental Setup
To evaluate POLAR , we conducted baseline experiments on three polarization detection tasks:
  • Task 1: Binary classification - determining whether a text is polarized.
  • Task 2: Multi-label classification - identifying polarization types (e.g., political, religious, ethnic).
  • Task 3: Multi-label classification - detecting polarization manifestations (e.g., incivility, stereotyping, dehumanization).
For data split, we used 70% for training, 10% for validation, and 20% for testing, as summarized in Table 1. We pursued two main experimental paradigms:
  1. 1.
    Fine-tuning Multilingual Language Models (MLMs): We fine-tuned six multilingual models including InfoXLM (Chi et al., 2021), LaBSE (Feng et al., 2022), RemBERT (Chung et al., 2021), XLM-R (Conneau et al., 2020), mBERT (Devlin et al., 2019), and mDeBERTa (He et al., 2023) for monolingual and crosslingual experiments.
  2. 2.
    Evaluating Large Language Models (LLMs) in Zero- and Few-shot Settings: We tested models including QWEN3-8B, LLAMA-3.1-8B, MIXTRAL-8X7B, and GPT4o/mini.
4.2 Results and Analysis
Monolingual Cosslingual
Task Lang. InfoXLM LaBSE RemBERT XLM-R mBERT mDeBERTa InfoXLM LaBSE RemBERT XLM-R mBERT mDeBERTa
1. Polarization Amharic 77.96 78.92 81.93 80.64 53.71 74.98 23.26 49.44 5.63 0.00 0.00 83.21
Arabic 58.39 61.94 67.19 70.42 52.48 54.68 29.68 35.87 27.76 2.78 37.79 25.63
English 72.09 69.84 75.43 76.08 72.77 74.94 1.90 51.68 53.41 46.89 33.44 53.45
German 29.75 63.46 67.50 67.78 58.45 64.38 0.62 51.97 64.59 41.87 46.05 25.94
Hausa 59.57 66.41 67.43 66.41 59.32 65.62 8.21 22.92 6.47 1.50 6.25 25.13
Spanish 38.69 64.04 70.98 57.00 59.31 54.16 45.70 68.66 67.26 21.09 68.34 62.17
Urdu 1.07 66.35 79.95 43.91 65.63 60.68 1.60 30.36 0.00 0.00 3.68 68.67
2. Types Amharic 24.22 38.13 43.65 25.71 17.93 25.00 11.56 20.65 2.49 0.00 0.00 26.76
Arabic 22.97 40.23 42.12 37.58 27.53 35.22 11.49 23.73 8.78 2.11 12.47 7.93
English 16.04 23.25 31.38 24.10 21.07 17.70 10.84 19.69 16.42 12.22 11.70 3.93
German 13.52 58.58 61.19 58.56 54.13 40.64 12.03 35.86 29.59 9.72 23.98 11.88
Hausa 18.56 19.14 17.38 18.09 19.61 18.87 3.20 9.97 3.91 1.39 5.78 5.90
Spanish 43.26 66.07 67.76 58.07 57.94 43.40 7.89 47.06 15.38 0.43 32.02 14.60
Urdu 27.14 51.94 51.60 45.38 38.02 33.13 3.77 13.62 6.33 0.00 3.94 20.30
3. Manifestations Amharic 43.52 47.56 47.63 43.17 33.18 43.29 15.57 27.07 8.64 0.00 0.00 43.58
Arabic 40.05 51.55 52.52 55.61 42.18 47.73 16.56 30.68 22.14 0.00 19.57 17.24
English 14.40 15.01 19.39 18.61 18.60 15.15 7.16 10.16 10.05 5.62 10.69 8.85
German 38.49 49.88 52.74 51.70 46.85 51.91 2.38 36.12 27.05 0.00 23.38 12.93
Hausa 19.23 20.04 19.18 18.89 18.74 18.93 5.74 5.86 3.77 3.12 6.19 6.43
Spanish 38.94 50.00 51.04 45.02 45.17 35.09 2.34 40.63 11.83 0.40 35.99 23.56
Urdu 34.26 52.20 53.64 41.32 45.90 47.01 2.10 19.16 11.54 0.00 1.88 48.58
Table 3: Average F1-Macro for all three tasks. In the monolingual settings, we train and evaluate the model for each language separately. In the crosslingual setting, we train on all languages within a language family (AfroAsian: Amharic, Arabic, Hausa, and Urdu), Germanic: English, German, and Spanish) except the target language, and evaluate on the test set of the target language. The best performance scores are highlighted in blue and orange, respectively.
4.2.1 Monolingual Setup Results
To benchmark the performance of state-of-the-art multilingual language models (MLMs) for polarization detection, we fine-tuned six prominent pretrained models: InfoXLM (Chi et al., 2021), LaBSE (Feng et al., 2022), RemBERT (Chung et al., 2021), XLM-R (Conneau et al., 2020), mBERT (Devlin et al., 2019), and mDeBERTa (He et al., 2023). Each model was evaluated in a monolingual (within-language) setting across three tasks: (i) binary polarization classification, (ii) multi-label polarization types classification, and (iii) multi-label polarization manifestations detection.
For Task 1, binary polarization classification, RemBERT consistently achieved the highest macro F1-scores for most languages, while XLM-R performed best in Arabic, English, and German. In the multi-label polarization types classification (Task 2), RemBERT again provided the top scores for Spanish, except for Hausa and Urdu, where mBERT and LaBSE performed best, respectively. For Task 3, multi-label polarization manifestations classification, the results were generally lower across all models, with RemBERT delivering the highest scores for most languages, except for Arabic and Hausa.
4.2.2 Cross-Lingual MLM Results
For cross-lingual transfer experiments, each MLM was trained on all other languages within the same language family (Afro-Asian or Germanic), excluding the target language, and evaluated on the target language’s test set for the same three tasks.
In the cross-lingual setup for polarization classification (Task 1), mDeBERTa achieved the highest F1-macro for Amharic, English, Hausa, and Urdu. For Arabic, German, and Spanish, mBERT, RemBERT, and LaBSE performed slightly better than the other models, respectively.
For Task 2, polarization types classification, performance was generally lower as expected in the cross-lingual scenario. LaBSE yielded the best results for four languages, while mDeBERTa performed better for Amharic and Urdu.
Similarly, for Task 3, polarization manifestation classification, the results were quite low overall, but LaBSE and mDeBERTa showed relatively better performance compared to the other models.
This suggests that general polarization detection transfers more easily across languages than the more nuanced identification of polarization types and manifestations, which are culturally and linguistically specific. Effective cross-lingual transfer likely requires aligned annotation schemes and fine-tuning strategies sensitive to these differences.
4.2.3 Zero and Few-shot LLM Performance
We evaluated five state-of-the-art LLMs (QWEN3-8B, LLAMA-3.1-8B, MIXTRAL-8X7B, GPT4o/mini) in a zero- and few-shot setting across languages and for the first ask. Larger models such as GPT4o and GPT4o-mini outperformed smaller ones, especially in capturing subtle polarization cues.
GPT 4o GPT 4o-mini Mistral-7B Llama-3.1-8B Qwen3-8B
Zero-shot Few-shot Zero-shot Few-shot Zero-shot Zero-shot Zero-shot
Amharic 71.30 68.74 55.40 62.59 71.92 36.32 58.09
Arabic 71.51 79.65 58.21 75.95 33.65 34.82 54.54
English 75.46 79.56 64.78 80.93 50.34 56.75 72.07
German 72.04 71.12 63.03 72.44 52.63 56.22 61.67
Hausa 34.73 44.87 28.21 44.26 23.42 18.51 25.06
Spanish 72.69 65.76 62.51 69.00 56.31 41.75 61.11
Urdu 72.25 78.83 59.19 74.66 70.47 67.26 69.63
Table 4: F1-Macro resulting from the zero- and few-shot LLM experiments with the POLAR  dataset. The highest value per language is highlighted in blue.
5 Discussion
Our multilingual and cross-cultural investigation into online polarization reveals that polarization is a deeply contextual and event-driven phenomenon. Its prevalence and intensity are closely tied to the sociopolitical climate, public discourse, and digital environment of each region. In societies experiencing major crises, such as the war in Ethiopia, religious tensions in Pakistan, and international conflicts like the wars in Ukraine and Gaza that influence the US election, we find polarization to be especially amplified on online platforms. Conversely, broader topical selection, moderation, or more diffuse societal tensions are associated with noticeably lower polarization.
The types of polarization evident in online discourse are shaped by the most salient local issues: political (Ethiopia, US, German), religious (Pakistan), ethnic (Ethiopia), or gender-based (Spain) divides tend to dominate where these subjects are at the forefront of public debate or conflict. This underscores the importance of accounting for each region’s unique sociopolitical history and cultural landscape when analyzing or intervening in digital polarization.
Polarization also manifests in varied rhetorical forms, including stereotyping, vilification, and exclusionary or hostile language. These patterns are particularly pronounced amid contentious issues and persistently divided communities, such as civil war, election, migration issues, religious divide, and so on.
From a computational perspective, our results demonstrate both the feasibility and the challenges of detecting polarization in multilingual settings. Binary polarization classification is relatively tractable for well-resourced languages and clear-cut contexts, especially when leveraging state-of-the-art multilingual models (Sec. 4. However, classifying the specific types and manifestations of polarization introduces significantly greater ambiguity and requires a much deeper, context-sensitive understanding. Performance on these more nuanced tasks remains limited. This highlights important directions for future work, including the integration of cultural signals and context or event understanding, the development of more robust multilingual embeddings, and the pursuit of consistent annotation (with a great deal of annotator training and follow up, specially for the crowdsourcing annotations). The wide variability across languages, contexts, and events further affirms the necessity of designing language, culture, and event specific approaches, particularly for underrepresented regions in global NLP research.
Taken together, these findings highlight that effective detection and mitigation of online polarization must move beyond generic or monocultural solutions. There is a pressing need for context-aware, adaptable methodologies that acknowledge both the universal and the local characteristics of polarized discourse, ensuring relevance and effectiveness across the world’s diverse digital landscapes.
6 Conclusion
In this study, we introduced POLAR, the first multilingual, multicultural, and multievent dataset for benchmarking online polarization. Our findings showed that polarization was deeply context-dependent, manifesting in diverse types and rhetorical forms shaped by local sociopolitical dynamics. Theoretically, we provide new evidence that polarization is a multifaceted, cross-lingual phenomenon strongly influenced by language, culture, and current events, rather than being uniform across settings. On a practical level, our benchmark revealed that while binary detection of polarization was feasible with current models, accurately identifying specific types and manifestations remained a persistent challenge, especially for multilingual and low-resource contexts.
This work established a foundation for further computational studies of digital polarization and for developing more culturally robust moderation and intervention tools. Looking ahead, future research should integrate richer contextual and cultural signals into model architectures, refine annotation guidelines for cross-regional consistency, and expand evaluations to include additional languages, social domains, and further events. By releasing our data and benchmarks, we aimed to catalyze further innovation toward nuanced detection and effective mitigation of online polarization worldwide.
Limitations
While POLAR represents an important step toward multilingual, multicultural, and multievent polarization analysis, several limitations remain. First, annotator understanding - particularly in crowdsourced setups - was sometimes limited, potentially impacting label quality. We mitigated this through strict quality assurance methods, including control questions, pre-study surveys, and ongoing annotator assessment, but some variability in interpretation may persist.
Second, in-house annotation, while yielding higher consistency, sometimes introduced psychological challenges for annotators given the sensitive or hostile nature of polarized content. To address this, we provided detailed instructions and support resources to reduce stress and clarify expectations, but some emotional burden may have remained.
Third, our choice of models is not exhaustive. Although we included several leading multilingual models and both open and closed LLMs. Adding more language-specific models in the future could improve results, especially for monolingual scenarios.
Finally, for some of the languages in our benchmark, the available data size is still limited, which may constrain the generalizability of model training and evaluation for those cases. Future work should expand dataset size and diversity, and explore language- or region-specific model development to better support underrepresented contexts.
Ethics Statement
This research uses only publicly available, anonymized data and addresses sensitive topics around polarization in diverse cultures. All annotation was conducted by native speakers using culturally appropriate guidelines; annotators were informed of the project’s social good aims, possible distress, and could opt out anytime. Annotators received prompt and fair compensation above local wage standards or per Prolific’s requirements. Despite rigorous protocols, labeling polarization remains subjective; we encourage responsible, ethically grounded use of this resource and discourage misuse.

### v2

@@
 POLAR: A Benchmark for Multilingual, Multicultural, and Multi-Event Online Polarization
-Usman Naseem1, Juan Ren1, Saba Anwar2, Sarah Kohail6, Rudy Alexandro Garrido Veliz2, Robert Geislinger2, Aisha Jabr6, Idris Abdulmumin5, Laiba Qureshi2, Aarushi Ajay Borkar2, Maryam Ibrahim Mukhtar7, Abinew Ali Ayele2,3, Ibrahim Said Ahmad7,8, Adem Ali2,3, Martin Semmann2, Shamsuddeen Hassan Muhammad4,7, Seid Muhie Yimam2 1Macquarie University, 2University of Hamburg, 3Bahir Dar University,4Imperial College London 5University of Pretoria, 6Zayed University, 7Bayero University Kano, 8Northeastern University
-Figure 1: Pipeline for POLAR construction: multi-platform data curation in 7 languages, annotation workflow with quality control, and benchmarking.
+Usman Naseem1, Juan Ren1, Saba Anwar2, Sarah Kohail6, Rudy Alexandro Garrido Veliz2, Robert Geislinger2, Aisha Jabr6, Idris Abdulmumin5, Laiba Qureshi2, Aarushi Ajay Borkar2, Maryam Ibrahim Mukhtar7, Abinew Ali Ayele2,3, Ibrahim Said Ahmad7,8, Adem Ali2,3, Martin Semmann2, Shamsuddeen Hassan Muhammad4,7, Seid Muhie Yimam2 1Macquarie University, 2University of Hamburg, 3Bahir Dar University, 4Imperial College London, 5University of Pretoria, 6Zayed University, 7Bayero University Kano, 8Northeastern University
 1 Introduction
-Online polarization, defined as sharp division and antagonism between social, political, or identity groups, has become a pervasive threat to democratic institutions, civil discourse, and social cohesion worldwide (Waller and Anderson, 2021; Iandoli et al., 2021).
-It is often fueled by biased or inflammatory content on social media, reinforcing echo chambers and undermining mutual understanding (Garimella, 2018).
+Online polarization, defined as sharp division and antagonism between social, political, or identity groups, has become a pervasive threat to democratic institutions, civil discourse, and social cohesion worldwide (waller2021quantifying; iandoli2021impact).
+It is often fueled by biased or inflammatory content on social media, reinforcing echo chambers and undermining mutual understanding (garimella2018polarization).
 Polarized discourse not only amplifies ideological divides but can also escalate into hate speech, harassment, and real-world violence.
 As such, early detection of polarization is critical to designing interventions that promote healthier online ecosystems.
+Figure 1: Pipeline for POLAR construction: multi-platform data curation in 7 languages, annotation workflow with quality control, and benchmarking.
 Despite growing attention, computational approaches to polarization suffer from major limitations.
@@
   3. 3.
-    Manifestation Identification: How is polarization rhetorically expressed (e.g., stereotyping, deindividuation, extreme language)?
+    Manifestation Identification: How is polarization rhetorically expressed (e.g., stereotyping/deindividuation, vilification, dehumanization, extreme language)?
 We design a cross-cultural annotation protocol tailored for each language’s sociopolitical context.
-Extensive qualitative analysis reveals how event salience, linguistic norms, and platform affordances shape polarization dynamics across languages.
 The complete pipeline can be found in Figure 1.
@@
 2 Related Work
-Online polarization has long been recognized as a threat to democracy and social cohesion, intensifying through social media echo chambers and biased content (Waller and Anderson, 2021; Iandoli et al., 2021; Garimella, 2018).
+Online polarization has long been recognized as a threat to democracy and social cohesion, intensifying through social media echo chambers and biased content (waller2021quantifying; iandoli2021impact; garimella2018polarization).
 As social media and other online platforms become key arenas for political and cultural discourse, the need for early detection and nuanced understanding of polarization has grown significantly.
 Such efforts are critical not only for content moderation, but also for peacebuilding, policy development, and responsible digital governance.
-Foundational research has defined polarization as both intergroup hostility and ingroup cohesion (Arora et al., 2022), and has highlighted its relationship with hate speech, fragmentation, and incivility (Bozdag and van den Hoven, 2020; Mathew et al., 2021).
-A growing body of research has documented the role of online spaces in intensifying polarization across different regions (Kubin and von Sikorski, 2021; Barberá, 2020; Gitlin, 2016; Soares and Recuero, 2021).
-However, most computational work focuses on high-resource languages and event- or region-specific datasets, limiting generalizability (Kubin and von Sikorski, 2021).
+Foundational research has defined polarization as both intergroup hostility and ingroup cohesion (arora2022polarization), and has highlighted its relationship with hate speech, fragmentation, and incivility (Mathew2021HateXplain).
+A growing body of research has documented the role of online spaces in intensifying polarization across different regions (kubin2021role; barbera2020social; gitlin2016outrage; soares2021hashtag).
+However, most computational work focuses on high-resource languages and event- or region-specific datasets, limiting generalizability (kubin2021role).
 This leaves a significant gap in our ability to generalize findings across cultures, languages, and events, especially in the Global South or multilingual regions.
 The lack of standardized datasets across languages has hindered progress in developing and evaluating polarization detection models with cross-lingual or cross-cultural capabilities.
-Recent shared tasks on hate speech and toxicity (Basile and others, 2019; Mohammad and others, 2021; Pamungkas et al., 2020; Ousidhoum et al., 2024) have expanded the language and domain coverage, yet remain less fine-grained regarding polarization’s diverse types and rhetorical manifestations.
+Recent shared tasks on hate speech and toxicity (basile2019semeval; pamungkas2020misogynistic; ousidhoum2024semeval) have expanded the language and domain coverage, yet remain less fine-grained regarding polarization’s diverse types and rhetorical manifestations.
 Our work addresses this gap by presenting the first comprehensive, fine-grained dataset benchmark for multilingual, multicultural, and multievent online polarization, enabling robust cross-lingual and context-aware modeling.
 3 POLAR Dataset Construction
+We aim to construct a multilingual, multi-event, and multi-platform social media dataset spanning multiple continents.
+The dataset is designed to cover Indo-European and Afro-Asiatic language families, with a particular focus on balancing linguistic diversity and resource availability.
+It includes high-resource languages such as English, Spanish, and German; medium-resource languages like Arabic and Urdu; and low-resource languages such as Hausa and Amharic.
+While several Indo-European languages are represented (e.g., English, Spanish, German, Urdu), the dataset also incorporates key Afro-Asiatic languages (e.g., Hausa, Arabic, Amharic) to ensure broader geographical and linguistic representation.
+The following section will discuss the dataset construction in detail.
+3.1 Operational Definitions
+In this study, polarization refers to the process or phenomenon in which opinions, beliefs, or behaviors become more extreme or divided, leading to a greater distance or conflict between differing groups.
+We also classify the types of polarization as follows:
+  • Political/ideological polarization: This type of extremism focuses on division, intolerance, and conflict between political parties and followers.
+  • Racial or ethnic polarization: This type of polarization focuses on ethnic identity or racial origin and incites division, intolerance, and conflict between ethnic groups or races.
+  • Religious polarization: This type of polarization focuses on religious identity and incites division, intolerance, and conflict between religious followers
+  • Gender polarization: This type of polarization refers to the exclusion, discrimination, and marginalization of individuals based on their gender.
+  • Sexual orientation polarization: This refers to the increasing division and distinction between different sexual orientations within society, often leading to heightened tensions, misunderstandings, conflicts, or marginalization among various groups.
+  • Other: Polarized texts targeting other groups/identities such as economy class, technology, media, polarization, etc are classified as others.
+Furthermore, we define a polarized text manifestations or tactics if it exhibits one or more of the following characteristics:
+  • Stereotype: A generalized belief that attributes specific characteristics to all members of a group, often neglecting individual differences, thereby reducing complex personalities to simplistic and uniform representations.
+  • Vilification: The act of defaming or demonizing a particular group, person, or entity by inciting fear, often through exaggeration, misrepresentation, or biased framing that portrays the subject negatively and harmfully.
+  • Dehumanization: The process of depriving a group or individual of their human qualities or personality by comparing them to animals, machines, or objects, or otherwise denying their humanity, dignity, or individuality.
+  • Extreme Language and Absolutism: The use of language that is extreme or makes definitive, all-encompassing statements, often involving words like "always," "never," "worst," or "best," and presenting issues in a dichotomous manner such as "us versus them" or "right versus wrong.
+  • Lack of Empathy or Understanding: The absence of compassion or recognition for other viewpoints or experiences in the text.
+  • Invalidation: The act of denying or dismissing the identity and existence of individuals or groups, thereby rejecting their sense of self and their presence.
+See Appendix A for further details and examples for each manifestation of polarization.
 3.2 Data Collection
@@
 Human experts curated keyword lists to reflect culturally and politically significant discourse across regions and events.
-Table 1 shows the languages covered, data splits and total number of instances annotated for each language.
+Table˜1 illustrates the compositions, size, and event coverage of the dataset.
 The Amharic dataset focuses on the Northern Ethiopia or Tigray War in Ethiopia; Arabic texts cover a broad array of social and economic topics; English data centers on US elections and international conflicts; the German corpus features election discourse and migration debates; the Hausa subset captures religious and ideological discussions in West and Central Africa; Spanish content spans abortion, migration, and indigenous/gender rights; and Urdu texts reflect political and sectarian divides.
-Language Source(s) Train Dev Test Total
-Amharic Facebook, X 3500 500 1000 5000
-Arabic Facebook, X, Threads, News 1482 212 424 2118
-English X, BlueSky, News 2117 303 605 3025
-German X, BlueSky, Reddit 2426 347 694 3467
-Hausa Facebook , X 3893 557 1113 5563
-Spanish X, BlueSky 1400 201 401 2002
-Urdu X 1960 280 560 2800
-Table 1: Dataset sources and split sizes for all three tasks as per each language.
+Language Source(s) Event Train Dev Test Total Instances
+Amharic Facebook, X Tigray War in Northern Ethiopia 3500 500 1000 5000
+Arabic Facebook, X, Threads, News Social and economic issues 1482 212 424 2118
+English X, BlueSky, News US elections and international conflicts 2117 303 605 3025
+German X, BlueSky, Reddit Election discourse and migration debates 2426 347 694 3467
+Hausa Facebook, X Religious and ideological discussions in Africa 3893 557 1113 5563
+Spanish X, BlueSky Abortion, migration, and indigenous/gender rights 1400 201 401 2002
+Urdu X Political and sectarian conflicts 1960 280 560 2800
+Table 1: Summary of dataset composition: sources, sizes, and event coverage by task and language.
 3.3 Annotation Process
 Given the cultural and linguistic breadth of POLAR , we developed detailed, multilingual annotation guidelines and deployed a hybrid strategy combining crowdsourcing with trained community annotators.
-Annotation Guidelines: We developed the guidelines in English and Amharic, and then translated and culturally adapted them for each target language. Annotators were instructed to:
+For crowdsourced annotators, we first filtered candidates based on their previous annotation experience and annotation agreement.
+Specifically, we selected those who had completed prior tasks with an agreement score of at least 0.8.
+We also conducted a pilot task consisting of 50 items to further filter annotators based on their agreement scores and average labeling time.
+For community annotators, we provided a training session followed by a pilot task to assess their performance before assigning the next batch of annotation tasks.
+During the annotation process, we continuously monitored annotator performance in a batch-wise manner to ensure that the overall process remained on track.
+Annotation Guidelines: We developed the guidelines in English and Amharic, and then translated and culturally adapted them for each target language.
+The English annotation guidelines are provided in Appendix A.
+Annotators were instructed to:
   • Identify whether a text is polarized;
   • If the text is classified as polarized, tag the type of polarization (political, racial/ethnic, religious, gender/sexual identity, other);
-  • If the text is classified as polarized, tag its manifestations (stereotyping, vilification, dehumanization, deindividuation, extreme language, lack of empathy, invalidation).
+  • If the text is classified as polarized, tag its manifestations (stereotyping/deindividuation, vilification, dehumanization, extreme language, lack of empathy, invalidation).
 Multiple labels were allowed due to the conceptual and contextual overlap often observed in polarized content.
-Annotator Recruitment and Annotation Process
-We used both expert and crowd-sourced annotation strategies.
+Annotator Recruitment and Annotation Process: We used both expert and crowd-sourced annotation strategies.
 Trained annotators used POTATO and Label Studio, while crowdsourced annotators used Mechanical Turk and Prolific.
@@
 Since different teams were responsible for dataset creation across various languages, we provide a detailed description of the annotation process for each language:
-Language Total Polarization Polarization Types Polarization Manifestations
-gender/sexual political religious racial/ethnic other vilification extreme_language stereotype invalidation lack_of_empathy dehumanization
-Amharic 5000 3753 29 3342 99 1297 189 2398 1527 2729 799 880 657
-Arabic 2118 338 102 178 86 171 220 413 341 283 240 175 102
-English 3025 1021 18 892 44 98 5 342 179 138 85 45 39
-German 3467 1596 241 1981 448 782 653 1371 1001 1420 1258 1124 568
-Hausa 5563 625 46 285 149 183 22 74 179 253 13 53 202
-Spanish 2002 1057 254 742 571 521 405 798 622 843 205 599 307
-Urdu 2800 1860 236 1533 557 423 126 1548 1409 979 670 632 501
-Table 2: Number of samples labeled positive for each annotation task across languages.
-Labels are grouped by task: Polarization (Task 1), Polarization Types (Task 2), and Polarization Manifestations (Task 3).
-  • Amharic: We used train annotators to annotated 5,000 samples in five batches using the POTATO annotation tool.
+  • Amharic: We used trained annotators to annotated 5,000 samples in five batches using the POTATO annotation tool.
 Annotators completed three rounds of guideline training and received continuous feedback, with ongoing supervision to ensure quality control.
@@
 Figure 3: Percentage of polarization per language
-(a) Types
-(b) Manifestation
-Figure 4: Annotation statistics for polarization classification, types, and manifestations
 3.4 Dataset Characteristics
-Polarization Prevalence: As shown in Figure 3, Amharic (75%) and Urdu (66%) has the highest polarization rates, tied to their sources - Northern Ethiopia’s Tigray conflict and Pakistan’s political and sectarian climate.
+Language Total Polarization Polarization Types Polarization Manifestations
+gender/sexual political religious racial/ethnic other vilification extreme_language stereotype invalidation lack_of_empathy dehumanization
+Amharic 5000 3753 29 3342 99 1297 189 2398 1527 2729 799 880 657
+Arabic 2118 338 102 178 86 171 220 413 341 283 240 175 102
+English 3025 1021 18 892 44 98 5 342 179 138 85 45 39
+German 3467 1596 241 1981 448 782 653 1371 1001 1420 1258 1124 568
+Hausa 5563 625 46 285 149 183 22 74 179 253 13 53 202
+Spanish 2002 1057 254 742 571 521 405 798 622 843 205 599 307
+Urdu 2800 1860 236 1533 557 423 126 1548 1409 979 670 632 501
+Table 2: Number of samples labeled positive for each annotation task across languages.
+Labels are grouped by task: Polarization (Task 1), Polarization Types (Task 2), and Polarization Manifestations (Task 3).
+We now analyze key trends and insights across the dataset based on annotated labels. Table 2 provides a detailed quantitative breakdown of positive labels for each annotation task.
+Polarization Prevalence: As shown in Figure 3 and further detailed in Table 2, Amharic (75%) and Urdu (66%) have the highest polarization rates, tied to their sources - Northern Ethiopia’s Tigray conflict and Pakistan’s political and sectarian climate.
 Spanish (53%), German (46%), and English (34%) fall in the mid-range.
@@
 Topic scope, moderation, and platform dynamics significantly influence polarization prevalence.
-Types of Polarization: Figure 4 show types of polarization.
-Political polarization dominates in Amharic (67%), German (57%), Urdu (55%), and Spanish (37%).
+Types of Polarization: As illustrated in Figure LABEL:fig:types and Table 2, political polarization dominates in Amharic (67%), German (57%), Urdu (55%), and Spanish (37%).
 Racial/ethnic and religious types also appear prominently in Spanish, Amharic, and German.
@@
 Spanish and Urdu show diverse forms of identity-based polarization, while political conflict dominates Amharic, Urdu, and German.
-Polarization Manifestations: Figure 4 show types of Manifestation.
-Stereotyping and vilification are most frequent in Urdu, Amharic, German, and Spanish.
+Polarization Manifestations: Figure LABEL:fig:manifestation and Table 2 show the types of manifestations.
+Stereotyping/deindividuation and vilification are most frequent in Urdu, Amharic, German, and Spanish.
 Urdu’s vilification rate is highest (55%).
@@
 Together, these analyses highlight how linguistic, cultural, and contextual variables - alongside event salience and platform dynamics - shape the structure and tone of online polarization.
-The POLAR dataset (see Table 2 for detailed statistics of POLAR dataset) offers a robust benchmark for examining these effects at scale.
+The POLAR dataset offers a robust benchmark for examining these effects at scale.
 4 Experimentation and Results
@@
   • Task 2: Multi-label classification - identifying polarization types (e.g., political, religious, ethnic).
-  • Task 3: Multi-label classification - detecting polarization manifestations (e.g., incivility, stereotyping, dehumanization).
-For data split, we used 70% for training, 10% for validation, and 20% for testing, as summarized in Table 1. We pursued two main experimental paradigms:
+  • Task 3: Multi-label classification - detecting polarization manifestations (e.g., stereotyping, vilification, dehumanization).
+For data split, we used 70% for training, 10% for validation, and 20% for testing, as summarized in Table˜1. We pursued two main experimental paradigms:
   1. 1.
-    Fine-tuning Multilingual Language Models (MLMs): We fine-tuned six multilingual models including InfoXLM (Chi et al., 2021), LaBSE (Feng et al., 2022), RemBERT (Chung et al., 2021), XLM-R (Conneau et al., 2020), mBERT (Devlin et al., 2019), and mDeBERTa (He et al., 2023) for monolingual and crosslingual experiments.
+    Fine-tuning Multilingual Language Models (MLMs): We fine-tuned six multilingual models including InfoXLM (chi2021infosxlm), LaBSE (feng2022language), RemBERT (chung2020rethinkingembeddingcouplingpretrained), XLM-R (conneau2019unsupervised), mBERT (devlin2019bert), and mDeBERTa (he2021debertav3) for monolingual and crosslingual experiments.
   2. 2.
-    Evaluating Large Language Models (LLMs) in Zero- and Few-shot Settings: We tested models including QWEN3-8B, LLAMA-3.1-8B, MIXTRAL-8X7B, and GPT4o/mini.
+[…]
 4.2 Results and Analysis
@@
 4.2.1 Monolingual Setup Results
-To benchmark the performance of state-of-the-art multilingual language models (MLMs) for polarization detection, we fine-tuned six prominent pretrained models: InfoXLM (Chi et al., 2021), LaBSE (Feng et al., 2022), RemBERT (Chung et al., 2021), XLM-R (Conneau et al., 2020), mBERT (Devlin et al., 2019), and mDeBERTa (He et al., 2023).
+To benchmark the performance of state-of-the-art multilingual language models (MLMs) for polarization detection, we fine-tuned six prominent pretrained models: InfoXLM (chi2021infosxlm), LaBSE (feng2022language), RemBERT (chung2020rethinkingembeddingcouplingpretrained), XLM-R (conneau2019unsupervised), mBERT (devlin2019bert), and mDeBERTa (he2021debertav3).
 Each model was evaluated in a monolingual (within-language) setting across three tasks: (i) binary polarization classification, (ii) multi-label polarization types classification, and (iii) multi-label polarization manifestations detection.
@@
 4.2.3 Zero and Few-shot LLM Performance
-We evaluated five state-of-the-art LLMs (QWEN3-8B, LLAMA-3.1-8B, MIXTRAL-8X7B, GPT4o/mini) in a zero- and few-shot setting across languages and for the first ask.
-Larger models such as GPT4o and GPT4o-mini outperformed smaller ones, especially in capturing subtle polarization cues.
-GPT 4o GPT 4o-mini Mistral-7B Llama-3.1-8B Qwen3-8B
-Zero-shot Few-shot Zero-shot Few-shot Zero-shot Zero-shot Zero-shot
-Amharic 71.30 68.74 55.40 62.59 71.92 36.32 58.09
-Arabic 71.51 79.65 58.21 75.95 33.65 34.82 54.54
-English 75.46 79.56 64.78 80.93 50.34 56.75 72.07
-German 72.04 71.12 63.03 72.44 52.63 56.22 61.67
-Hausa 34.73 44.87 28.21 44.26 23.42 18.51 25.06
-Spanish 72.69 65.76 62.51 69.00 56.31 41.75 61.11
-Urdu 72.25 78.83 59.19 74.66 70.47 67.26 69.63
+We evaluated seven state-of-the-art LLMs (Qwen3-8B, LLaMA-3.1-8B, Mixtral-8x7B, GPT-4o, GPT-4o-mini, GPT-4.1-mini, and Llama-4-Scout-17B-16E-Instruct) across languages.
+The prompts used for evaluation are provided in the Appendix B.
+We explicitly provided definitions of tasks and leveraged the models’ chain-of-thoughts capabilities by encouraging reasoning before labeling.
+In the few-shot setting, we provided 4 example instances for each task.
+To assess the impact of prompt language, we also compared model performance when prompted in English versus the respective target language.
+As Table  4 illustrates, the closed-source models GPT-4o, GPT-4o mini, and GPT-4.1 mini, consistently outperform the other four LLMs, demonstrating superior performance in polarization recognition.
+However, their performance on Hausa remains notably low, highlighting both the challenge of understanding low-resource languages and the value of our multilingual dataset in exposing the limitations.
+In contrast, smaller open-source models, such as Qwen3-8B, LLaMA-3.1-8B, and Mixtral-8x7B, achieve lower F1-Macro scores and perform inconsistently across languages.
+Their F1-Macro scores improve by 5 10% when provided with few-shot examples.
+However, the larger LLaMA model, LLaMA-4-scout-17B, demonstrates performance comparable to GPT-4o in polarization classification.
+An interesting phenomenon is observed: Smaller models show improved classification performance in the few-shot setting, while larger models such as the GPT series and LLaMA-4-17B do not consistently gain additional improvement with few-shot examples.
+This suggests that few-shot learning offers advantages to smaller models than to larger ones.
+Task Lang GPT 4o GPT 4o-mini GPT 4.1-mini Mistral-7B Qwen3-8B Llama-3.1-8B Llama4-17B
+Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot
+1. Polarization Amharic 71.30 68.74 55.40 58.04 63.52 63.20 44.84 48.24 55.60 65.51 50.98 58.84 69.96 67.91
+Arabic 71.51 79.65 58.21 68.41 76.20 79.58 33.65 53.06 65.95 70.76 61.51 67.21 73.58 77.45
+English 75.46 79.56 64.78 80.30 77.43 77.48 53.78 70.00 69.09 66.86 66.55 69.90 76.90 75.48
+German 72.04 71.12 63.03 68.49 73.77 69.60 56.17 61.77 65.42 68.76 59.90 65.40 70.89 70.35
+Hausa 34.73 44.87 28.21 34.07 39.45 48.06 19.82 28.15 45.91 49.18 38.73 40.80 47.85 54.53
+Spanish 72.69 65.76 62.51 67.96 71.28 69.22 51.24 61.00 65.42 68.76 60.20 62.53 66.11 66.27
+Urdu 72.25 78.83 59.19 69.92 73.15 72.63 48.27 54.79 66.99 70.89 55.61 59.77 73.93 68.50
+2. Types Amharic 39.67 36.59 33.8 34.89 32.49 34.44 44.21 63.24 63.07 75.90 78.26
+Arabic 35.00 39.30 37.18 36.82 33.07 36.77 49.39 53.91 53.22 52.86 55.03
+English 34.96 35.67 33.79 38.73 35.22 38.22 48.28 68.50 67.37 62.71 67.09
+German 58.86 58.58 54.36 57.94 52.70 55.15 44.10 71.82 70.62 73.89 76.20
+Hausa 18.96 19.11 16.90 18.15 17.00 18.86 43.34 67.24 70.47 71.75 74.66
+Spanish 48.98 52.93 52.82 55.09 48.89 49.76 43.15 65.96 67.87 69.82 71.43
+Urdu 55.6 59.64 53.97 54.78 51.82 55.09 48.40 63.36 64.83 60.01 61.52
+3. Manifestations Amharic 44.09 45.59 40.01 44.78 40.53 42.49 0 0 0 0 0
+Arabic 51.90 56.90 46.96 51.01 52.57 60.76 0 0 0 0 0
+English 22.95 25.87 20.43 25.78 28.41 26.67 0 0 0 0 0
+German 42.84 45.16 43.33 53.78 37.62 41.23 0 0 0 0 0
+Hausa 8.58 8.14 8.42 8.37 8.55 9.33 0 0 0 0 0
+Spanish 41.00 40.99 39.85 46.33 38.82 40.66 0 0 0 0 0
+Urdu 52.89 54.43 48.25 58.17 48.34 51.89 0 0 0 0 0
 Table 4: F1-Macro resulting from the zero- and few-shot LLM experiments with the POLAR  dataset. The highest value per language is highlighted in blue.
@@
 Conversely, broader topical selection, moderation, or more diffuse societal tensions are associated with noticeably lower polarization.
-The types of polarization evident in online discourse are shaped by the most salient local issues: political (Ethiopia, US, German), religious (Pakistan), ethnic (Ethiopia), or gender-based (Spain) divides tend to dominate where these subjects are at the forefront of public debate or conflict.
+The types of polarization evident in online discourse are shaped by the most salient local issues: political (e.g., in Ethiopia, the US, and Germany), religious (e.g., in Pakistan), ethnic (e.g., in Ethiopia), or gender-based (e.g., in Spain) divides tend to dominate where these subjects are at the forefront of public debate or conflict.
 This underscores the importance of accounting for each region’s unique sociopolitical history and cultural landscape when analyzing or intervening in digital polarization.
-Polarization also manifests in varied rhetorical forms, including stereotyping, vilification, and exclusionary or hostile language.
+Polarization also manifests in varied rhetorical forms, including stereotyping/deindividuation, vilification, and exclusionary or hostile language.
 These patterns are particularly pronounced amid contentious issues and persistently divided communities, such as civil war, election, migration issues, religious divide, and so on.
@@
 Binary polarization classification is relatively tractable for well-resourced languages and clear-cut contexts, especially when leveraging state-of-the-art multilingual models (Sec.
-4.
+4).
 However, classifying the specific types and manifestations of polarization introduces significantly greater ambiguity and requires a much deeper, context-sensitive understanding.
@@
 Despite rigorous protocols, labeling polarization remains subjective; we encourage responsible, ethically grounded use of this resource and discourage misuse.
+As Table 5 demonstrated, the inter-annotation agreement scores range from 0.2 to 0.6, corresponding to fair to moderate levels of agreement according to established benchmarks (landis1977kappa).
+This level of agreement further highlights the inherent difficulty of polarization recognition and underscores the necessity and value of our proposed work.
+Language Annotation Platform Kappa Score
+Amharic Potato 0.65
+Arabic Potato 0.296
+English MTurk 0.52
+German Prolific 0.29
+Hausa Potato 0.48
+Spanish MTurk 0.24
+Urdu Potato 0.55
+Table 5: Dataset annotation platforms and inter-annotator agreement scores
+Appendix A Annotation Guidelines
+Annotation Guidelines
+To minimize potential misunderstandings, we offer annotation guidelines customized for each event, accompanied by event-specific examples.
+These guidelines are in the languages of the target datsets.
+The English version is provided below.
+This guideline aims to assess whether social media messages reflect attitude polarization and to categorize the various types and manifestations of this polarization.
+The dataset consists of content sourced from platforms such as Facebook and Twitter.
+Warning to annotators: You may encounter polarized and hateful content during this task. Please take breaks as needed and prioritize your mental well-being.
+If you choose to participate, please read the following sections carefully. To exit the study at any time, click “Cancel” or “Quit” and provide feedback if possible.
+Task 1: Polarization Classification (Yes/No)
+Polarization refers to increasingly extreme, divided beliefs or behaviors between opposing groups. Attitude polarization includes:
+  • Negative attitudes toward out-groups
+  • Blind support for in-groups
+  • Stereotyping, vilification, dehumanization, or intolerance
+Texts should be labeled:
+  • Yes – if the message clearly reflects attitude polarization
+  • No – if it does not show any polarization indicators
+Note: Always consider the overall context and meaning, not just individual words.
+Task 2: Type of Polarization (if “Yes” in Task 1)
+If the message is polarized, select all applicable types:
+  1. 2.1
+    Political/Ideological Polarization – Conflict between political parties or ideologies
+  2. 2.2
+    Racial/Ethnic Polarization – Division based on race or ethnicity
+  3. 2.3
+    Religious Polarization – Conflict based on religious beliefs
+  4. 2.4
+    Gender Polarization – Discrimination based on gender
+  5. 2.5
+    Sexual Orientation Polarization – Tension or marginalization based on sexual orientation
+  6. 2.6
+    Other – e.g., based on economic class, technology, media, etc.
+Note: Multiple types may be selected if applicable.
+Task 3: Manifestations of Polarization (if “Yes” in Task 1)
+Classify the tactics used to express polarization in the message. A text is considered polarizing if it includes one or more of the following:
+  1. 3.1
+    Stereotyping – Generalizing traits to all group members Example: “Women are weak.”
+  2. 3.2
+    Vilification – Defaming or demonizing a group Example: “Migrants are traitors.”
+  3. 3.3
+    Dehumanization – Stripping away human qualities Example: “They are cockroaches.”
+  4. 3.4
+    Deindividuation – Ignoring individual identities Example: “All migrants are thieves.”
+  5. 3.5
+    Extreme Language or Absolutism – Using words like “always,” “never,” or “worst” Example: “We can never trust them.”
+  6. 3.6
+    Lack of Empathy or Understanding – Dismissing others’ perspectives Example: “Wearing the hijab is extremist.”
+  7. 3.7
+    Invalidation – Denying a group’s identity or existence Example: “There is no nation called Palestine.”
+Thank you for your valuable contribution to this research.
+Appendix B Prompt for Text Classification
+Task 1: Polarization Detection – Zero Shot
+Given the following texts in [X] language, classify them as polarized or non-polarized based on the following definition:
+Definition of Polarization: Polarization refers to the phenomenon where opinions, beliefs, or behaviors become more extreme or divided, resulting in increased separation and conflict between different groups.
+These groups may be based on politics or ideology, ethnicity or race, religion, gender, or sexual orientation.
+Polarization involves negative attitudes such as stereotyping, denigration, dehumanization, de-individualization, or intolerance toward individuals or groups outside one’s own, while exhibiting blind support and solidarity toward those within the group.
+Reasoning: Provide the reason why the text demonstrates polarization.
+Label: Return True if the text is polarized, otherwise False. Do not return any other value.
+Task 1: Polarization Detection – Few Shot
+Same instructions as above. Examples provided:
+  • Text: Black immigrants should leave the USA.
+    Reasoning: This statement stereotypes a group based on race and immigration status, implying a negative attitude and exclusion.
+    Label: True
+  • Text: Africans are lazy because they are backward.
+    Reasoning: Negative trait assignment based on ethnicity, showing polarization.
+    Label: True
+  • Text: Muslims are terrorists.
+    Reasoning: Promotes harmful stereotypes and dehumanization.
+    Label: True
+  • Text: African women are beautiful.
+    Reasoning: Positive generalization, not polarizing.
+    Label: False
+Task 2: Polarized Text Typing – Zero Shot
+Classify each polarized text into one or more of the following types:
+  • Gender/Sexual: Discrimination based on gender or sexual orientation
+  • Political: Division or conflict between political ideologies
+  • Religious: Conflict between religious groups
+  • Racial/Ethnic: Conflict or division based on race or ethnicity
+  • Other: Related to economy, technology, media, etc.
+Output: A list of five binary values in the order: [Gender/Sexual, Political, Religious, Racial/Ethnic, Other]
+Few Shot Examples
+  • Text: Black immigrants should leave the USA.
+    Output: [0, 0, 0, 1, 0]
+  • Text: Muslims are terrorists.
+    Output: [0, 0, 1, 0, 0]
+  • Text: Women should be confined to traditional roles.
+    Output: [1, 0, 0, 0, 0]
+  • Text: The economy is failing because of government policies.
+    Output: [0, 1, 0, 0, 0]
+Task 3: Polarization Manifestations – Zero Shot
+Analyze and identify the following manifestations in each text. Output a list of six binary values in this order:
+[Vilification, Extreme Language, Stereotype, Invalidation, Lack of Empathy, Dehumanization]
+Definitions
+  • Stereotype/Deindividuation: Generalizes traits, ignores individuality.
+  • Vilification: Defames a group or person, inciting fear.
+  • Dehumanization: Strips humanity, compares people to animals or objects.
+  • Extreme Language: Uses absolutes like always, never, us vs. them.
+  • Lack of Empathy: Dismisses others’ experiences or views.
+  • Invalidation: Denies the identity or existence of a group.
+Few Shot Examples
+  • Text: Black immigrants should leave the USA.
+    Output: [1, 1, 1, 1, 1, 0]
+  • Text: Africans are lazy because they are backward.
+    Output: [1, 1, 1, 1, 1, 0]
+  • Text: Muslims are terrorists.
+    Output: [1, 1, 1, 1, 1, 1]
+  • Text: African women are beautiful.
+    Output: [0, 0, 0, 0, 0, 0]
+Appendix C Dataset Samples
+Table 6: Examples of polarized statements with annotated types and manifestations.
+Types: P: Political, Re: Religious, Ra:Racial; Manifestations: V: Vilification, EL: Extreme Language, S: Stereotype, I: Invalidation, LE: Lack of Empathy, D: Dehumanization.
+Text T2 T3
+"Denazification" as an excuse to explain Russia’s deployment of troops to Ukraine, alleging that Hitler was of Jewish ancestry P V, EL
+"Heck, I voted for Biden three times." Signed, Liberal Asshats in Swing States P EL, S, D
+Satellite images confirm civilian massacres in towns during Russian military occupation P V
+"Sufficient evidence of voter fraud?" Where? marjorieisaninsurrectionist P EL, I
+"They" are now the deep state. P S, I, D
+15000 dead children claimed by Hamas who uses those children as human shields. Yes yes we covered that already. Now you’re just talking in circles P,Ra V, EL, I, LE, D
+5 Black Officers Awarded \$16M After White Colleague’s Racial Comments R V
+A Federal Investigation needs to be launched against TWITTER Fraud and Election Interference P EL, I
+A Jewish state was created because no one else wanted them. This is why. Zionism and Naziism are the same disease. P+Re+Ra V, EL, S, I, LE, D
+A lot of lying by the radical left. P V, S, I, D
+A rational and just society probably wouldn’t allow school shootings to be part of their National identity P EL, S, I, D
+A relationship between the Apartheid state and a tiny tyrant state. Free Palestine P V, I, LE, D
+A small price to pay for Ukrainian sovereignty and our green future. Stop whining. P I, LE
+A traitor today, a traitor tomorrow, a traitor always! P V
+A valid protest of a rigged election, that ended up having some sort of kerfuffle. P V, I
+A very moral army. IDF Gaza Israel P V, LE
+Text Lang Label
+Gibt sogar recht viele Ex-Muslime. Alles sehr vernünftige Leute. DE TRUE
+There are actually quite a few ex-Muslims. All very reasonable people. EN
+Kriegstreiber! Auch ihr, die Propagandapresse!!! DE TRUE
+Warmongers! You too, the propaganda press!!! EN
+Steuerzahler und Juden DE TRUE
+Taxpayers and Jews EN
+Ja die Türken leiden sehr stark unter ihrem links grün versifften Erdogan. Muss man wissen. :D DE TRUE
+Yes, the Turks are suffering greatly under their left-green filthy Erdogan. Good to know. :D EN
+Die Klima Armee Fraktion wird kommen. DE TRUE
+The Climate Army faction will come. EN
+Denk ich nicht, nein. Migration ist deren Kernthema. DE TRUE
+I don’t think so, no. Migration is their core issue. EN
+"Es gibt bereits vereinzelte Einsätze von Bodentruppen in Gaza. @URL" DE FALSE
+"There are already isolated deployments of ground troops in Gaza. @URL" EN
+Lass dich nicht aufhalten. Es gibt eine ukrainische Fremdenlegion. Da kannst du dich melden. DE FALSE
+Don’t let yourself be stopped. There is a Ukrainian foreign legion. You can sign up there. EN
+Klingt für mich nach Long COVID DE FALSE
+Sounds like long COVID to me. EN
+Wenn ich schon wieder dieses Woke lese. Der AfD ist sogar unser Grundgesetz zu Woke. Hört auf das Wort so inflationär zu gebrauchen. DE FALSE
+When I read this "woke" again. Even our constitution is too woke for the AfD. Stop using the word so inflationarily. EN
+CO2 ist gut für Pflanzen DE FALSE
+CO2 is good for plants. EN
+Table 7: German samples with labels and English translations
+Text Lang Label
+y vos un random judio que comenta pelotudeces para poder figurar en algun lado SPA TRUE
+And you, a random Jew who comments nonsense just to get noticed somewhere. EN
+quiero comer indio viejo SPA TRUE
+I want to eat old Indian. EN
+es que a dia de hoy parece que la diferencia es nazis con aborto y nazis sin aborto SPA TRUE
+Nowadays it seems that the difference is Nazis with abortion and Nazis without abortion. EN
+orgulloso de sacarle la conchetumare a un pastor evangelico por meterse en weas que no debia. SPA TRUE
+Proud to beat the hell out of an evangelical pastor for meddling in things he shouldn’t. EN
+el tribunal superior del partido corrupto SPA TRUE
+The supreme court of the corrupt party. EN
+claudia es judia, que no manche SPA FALSE
+Claudia is Jewish, don’t stain. EN
+la escuela feminista de pintura SPA FALSE
+The feminist school of painting. EN
+la cancion que me representa ahora que estoy deportado de los estados unidos. SPA FALSE
+The song that represents me now that I am deported from the United States. EN
+aunque, siendo justos, esa foto es de una deportacion de 2018, cuando trump estaba gobernando. SPA FALSE
+Although, to be fair, that photo is from a 2018 deportation, when Trump was in power. EN
+la propuesta de k!nadim me parece muy interesante de cara a eurovision, es facilona y divertida #benidormfestsemi1 SPA FALSE
+The proposal by k!nadim seems very interesting to me for Eurovision; it’s easy and fun. #benidormfestsemi1 EN
+Table 8: Spanish sentences with labels and English translations.

### v3

@@
-POLAR: A Benchmark for Multilingual, Multicultural, and Multi-Event Online Polarization
+and Multi-Event Online Polarization
 Abstract
 Online polarization poses a growing challenge for democratic discourse, yet most computational social science research remains monolingual, culturally narrow, or event-specific.
-We introduce POLAR, a multilingual, multicultural, and multievent dataset with over 23k instances in seven languages from diverse online platforms and real-world events.
-Polarization is annotated along three axes: presence, type, and manifestation, using a variety of annotation platforms adapted to each cultural context.
-We conduct two main experiments: (1) we fine-tune six multilingual pretrained language models in both monolingual and cross-lingual setups; and (2) we evaluate a range of open and closed large language models (LLMs) in few-shot and zero-shot scenarios.
-Results show that while most models perform well on binary polarization detection, they achieve substantially lower scores when predicting polarization types and manifestations.
-These findings highlight the complex, highly contextual nature of polarization and the need for robust, adaptable approaches in NLP and computational social science.
+We introduce POLAR, a multilingual, multicultural, and multi-event dataset with over 110K instances in 22 languages drawn from diverse online platforms and real-world events.
+Polarization is annotated along three axes, namely detection, type, and manifestation, using a variety of annotation platforms adapted to each cultural context.
+We conduct two main experiments: (1) fine-tuning six pretrained small language models; and (2) evaluating a range of open and closed large language models in few-shot and zero-shot settings.
+The results show that, while most models perform well in binary polarization detection, they achieve substantially lower performance when predicting polarization types and manifestations.
+These findings highlight the complex, highly contextual nature of polarization and demonstrate the need for robust, adaptable approaches in NLP and computational social science.
 All resources will be released to support further research and effective mitigation of digital polarization globally.
-POLAR: A Benchmark for Multilingual, Multicultural, and Multi-Event Online Polarization
-Usman Naseem1, Juan Ren1, Saba Anwar2, Sarah Kohail6, Rudy Alexandro Garrido Veliz2, Robert Geislinger2, Aisha Jabr6, Idris Abdulmumin5, Laiba Qureshi2, Aarushi Ajay Borkar2, Maryam Ibrahim Mukhtar7, Abinew Ali Ayele2,3, Ibrahim Said Ahmad7,8, Adem Ali2,3, Martin Semmann2, Shamsuddeen Hassan Muhammad4,7, Seid Muhie Yimam2 1Macquarie University, 2University of Hamburg, 3Bahir Dar University, 4Imperial College London, 5University of Pretoria, 6Zayed University, 7Bayero University Kano, 8Northeastern University
+POLAR: A Benchmark for Multilingual, Multicultural,
+and Multi-Event Online Polarization
+Usman Naseem1, Robert Geislinger2, Juan Ren1, Sarah Kohail3, Rudy Garrido Veliz2, P Sam Sahil2,4, Yiran Zhang1, Marco Antonio Stranisci6,7, Idris Abdulmumin5, Özge Alacam8, Cengiz Acartürk9, Aisha Jabr3,Saba Anwar2, Abinew Ali Ayele10, Simona Frenda7,11, Alessandra Teresa Cignarella7, Elena Tutubalina13,14,15, Oleg Rogov13,16,17, Aung Kyaw Htet1, Xintong Wang2, Surendrabikram Thapa18, Kritesh Rauniyar1, Tanmoy Chakraborty19, Arfeen Zeeshan19, Dheeraj Kodati20, Satya Keerthi21, Sahar Moradizeyveh1, Firoj Alam22,23, Arid Hasan24, Syed Ishtiaque Ahmed24, Ye Kyaw Thu25, Shantipriya Parida26, Ihsan Ayyub Qazi27, Lilian Wanzare28, Nelson Odhiambo Onyango28, Clemencia Siro29, Jane Wanjiru Kimani30, Ibrahim Said Ahmad31,32, Adem Chanie Ali2,10, Martin Semmann2, Chris Biemann2, Shamsuddeen Hassan Muhammad33, Seid Muhie Yimam2 1Macquarie University, 2University of Hamburg, 3Zayed University, 4HKBK College of Engineering, 5University of Pretoria, 6University of Turin, 7aequa-tech, 8Bielefeld University, 9Jagiellonian University, 10Bahir Dar University, 11Heriot-Watt University, 12Ghent University, 13AIRI, 14KFU, 15HSE University, 16MTUCI, 17Skoltech, 18Virginia Tech, 19ITT Delhi, 20ABV-IIITM, 21Mahindra University, 22Qatar Computing Research Institute, 23Hamad Bin Khalifa University, 24University of Toronto, 25LU Lab., Myanmar, 26AMD Silo AI, 27Lahore University of Management Sciences, 28Maseno University, 29Centrum Wiskunde & Informatica, 30Jomo Kenyatta University of Agriculture and Technology, 31Bayero University Kano, 32Northeastern University, 33Imperial College London,
 1 Introduction
-Online polarization, defined as sharp division and antagonism between social, political, or identity groups, has become a pervasive threat to democratic institutions, civil discourse, and social cohesion worldwide (waller2021quantifying; iandoli2021impact).
-It is often fueled by biased or inflammatory content on social media, reinforcing echo chambers and undermining mutual understanding (garimella2018polarization).
-Polarized discourse not only amplifies ideological divides but can also escalate into hate speech, harassment, and real-world violence.
-As such, early detection of polarization is critical to designing interventions that promote healthier online ecosystems.
-Figure 1: Pipeline for POLAR construction: multi-platform data curation in 7 languages, annotation workflow with quality control, and benchmarking.
-Despite growing attention, computational approaches to polarization suffer from major limitations.
-First, most existing datasets focus on English or high-resource languages, reflecting a widespread trend across NLP tasks that ignores the rich diversity of linguistic and sociocultural contexts in which polarization manifests.
-Second, current benchmarks are often event-specific or monodomain, such as U.S.
-elections or Western political debates, limiting their generalizability.
-Third, the conceptualization of polarization in NLP has largely been binary or topic-focused, overlooking the multifaceted ways in which polarization is expressed through vilification, dehumanization, stereotyping, or other rhetorical tactics.
-To address these gaps, we introduce POLAR a novel multilingual, multicultural, and multievent dataset for fine-grained polarization detection.
-It spans seven languages across diverse regions, including low-resource languages such as Amharic and Hausa.
-Our data is sourced from various platforms (e.g., Twitter/X, Facebook, BlueSky, Reddit, and local news outlets), reflecting authentic, event-driven discourse ranging from armed conflict (e.g., the Tigray War) to social justice movements (e.g., abortion rights, migration crises).
+Online polarization, defined as sharp division and antagonism between social, political, or identity groups, has become a pervasive threat to democratic institutions, civil discourse, and social cohesion worldwide (Waller and Anderson, 2021).
+It is often fueled by biased or inflammatory content in digital media, strengthening echo chambers and undermining mutual understanding (Garimella, 2018).
+Polarized discourse amplifies ideological divides and can escalate into hate speech, harassment, and real-world violence Piazza (2023); Martínez-España et al.
+(2024).
+Therefore, early detection of polarization is essential for designing interventions that promote healthier online ecosystems.
+Figure 1: Pipeline for POLAR construction. Data curation in 22 languages, annotation workflow, and benchmarking.
+Despite increasing attention, computational approaches to polarization face several limitations.
+First, most existing datasets focus on English or high-resource languages, reflecting a widespread trend across NLP tasks that ignores the rich diversity of linguistic and sociocultural contexts in which polarization manifests Simchon et al.
+(2022); Piazza (2023); Rojo Martínez (2025).
+Second, previous studies are event-specific or monodomain, such as U.S.
+elections or Western political debates, limiting their generalizability Demszky et al.
+(2019); Casal Bértoa and Rama (2021); Sinno et al.
+(2022); Piazza (2023).
+Third, the conceptualization of polarization in NLP has largely been binary or topic-focused Hofmann et al.
+(2022), overlooking the multifaceted ways in which polarization is expressed through vilification, dehumanization, stereotyping, or other rhetorical tactics Donohue and Hamilton (2022).
+These tactics are often employed in political rhetoric, social debates, or campaigns to solidify support within a group and increase hostility to others.
+To address these gaps, we introduce POLAR, a large-scale, multilingual, multicultural, and multievent dataset for fine-grained polarization detection.
+POLAR supports 22 languages spanning seven language families and balances high-, medium-, and low-resource languages (see Table˜5).
+The wide extent of our efforts can be seen in Figure 2.
 Unlike prior work, POLAR supports three complementary tasks:
-  1. 1.
-    Binary Polarization Detection: Is a message polarized or not?
-  2. 2.
-    Polarization Type Classification: What social dimension is targeted (e.g., political, religious, racial)?
-  3. 3.
-    Manifestation Identification: How is polarization rhetorically expressed (e.g., stereotyping/deindividuation, vilification, dehumanization, extreme language)?
-We design a cross-cultural annotation protocol tailored for each language’s sociopolitical context.
-The complete pipeline can be found in Figure 1.
-We benchmark a range of multilingual language models (MLMs) and large language models (LLMs) under zero-shot, few-shot, and cross-lingual scenarios.
+  • Binary Polarization Detection: Determine whether a given text expresses polarization. We refer to this task as PolarDetect.
+  • Polarization Type Classification: Identify the social dimension underlying the polarization (e.g., political, religious, racial). We refer to this task as PolarType.
+  • Manifestation Identification: Detect how polarization is rhetorically manifested, including strategies such as stereotyping, deindividuation, vilification, dehumanization, extreme language, and other rhetorical devices.
+We refer to this task as PolarManifest.
+For each task, we develop a cross-cultural annotation protocol tailored for each language’s sociopolitical context.
+The complete data construction pipeline is illustrated in Figure˜1.
+We benchmark a range of Small Language Models (SLMs) and Large Language Models (LLMs) under zero-shot and few-shot settings.
 Our experiments highlight the challenges of generalization and the limitations of current models in capturing nuanced rhetorical patterns across languages.
 Our contributions are as follows:
-  • We release POLAR, the first large-scale, multilingual, fine-grained dataset for polarization detection across 7 languages and diverse global events.
+  • We release POLAR, the first large-scale, multilingual, fine-grained dataset for polarization analysis across 22 languages and diverse global events, comprising 110K instances.
   • We define a taxonomy of polarization types and manifestations, operationalized through a robust cross-lingual annotation protocol.
-  • We provide comprehensive benchmarks using state-of-the-art MLMs and LLMs across multiple evaluation settings (monolingual, cross-lingual, and few-shot).
-Figure 2: Languages represented in the dataset, covering diverse linguistic and regional contexts.
+  • We provide comprehensive benchmarks using state-of-the-art SLMs and LLMs across multiple evaluation settings.
+Figure 2: Languages represented in a world map covered by POLAR, covering diverse linguistic and regional contexts.
+The language and societal context can present itself across varied areas.
+Language assignments to countries and regions are approximate.
 2 Related Work
-Online polarization has long been recognized as a threat to democracy and social cohesion, intensifying through social media echo chambers and biased content (waller2021quantifying; iandoli2021impact; garimella2018polarization).
+Online polarization poses a threat to social cohesion, exacerbated by social media echo chambers and biased content (Waller and Anderson, 2021; Iandoli et al., 2021; Garimella, 2018).
 As social media and other online platforms become key arenas for political and cultural discourse, the need for early detection and nuanced understanding of polarization has grown significantly.
-Such efforts are critical not only for content moderation, but also for peacebuilding, policy development, and responsible digital governance.
-Foundational research has defined polarization as both intergroup hostility and ingroup cohesion (arora2022polarization), and has highlighted its relationship with hate speech, fragmentation, and incivility (Mathew2021HateXplain).
-A growing body of research has documented the role of online spaces in intensifying polarization across different regions (kubin2021role; barbera2020social; gitlin2016outrage; soares2021hashtag).
-However, most computational work focuses on high-resource languages and event- or region-specific datasets, limiting generalizability (kubin2021role).
-This leaves a significant gap in our ability to generalize findings across cultures, languages, and events, especially in the Global South or multilingual regions.
+Polarization detection is important for content moderation, peace building, policy development, responsible digital governance and healthy democracy.
+Foundational research has defined polarization as both intergroup hostility and blind ingroup cohesion (Arora et al., 2022), and has highlighted its relationship with hate speech, fragmentation, and incivility (Mathew et al., 2021).
+A growing body of research has documented the role of online spaces in intensifying polarization across different regions (Kubin and von Sikorski, 2021; Barberá, 2020; Gitlin, 2016; Soares and Recuero, 2021).
+However, most computational work focuses on high-resource languages and event- or region-specific datasets, limiting generalizability (Kubin and von Sikorski, 2021).
+This leaves a significant gap in our ability to generalize findings across cultures, languages, and events, especially for low-resource languages or multilingual regions.
 The lack of standardized datasets across languages has hindered progress in developing and evaluating polarization detection models with cross-lingual or cross-cultural capabilities.
-Recent shared tasks on hate speech and toxicity (basile2019semeval; pamungkas2020misogynistic; ousidhoum2024semeval) have expanded the language and domain coverage, yet remain less fine-grained regarding polarization’s diverse types and rhetorical manifestations.
-Our work addresses this gap by presenting the first comprehensive, fine-grained dataset benchmark for multilingual, multicultural, and multievent online polarization, enabling robust cross-lingual and context-aware modeling.
+Recent shared tasks on hate speech and toxicity (Basile et al., 2019; Pamungkas et al., 2020) have expanded the language and domain coverage, yet remain less fine-grained regarding polarization’s diverse types and rhetorical manifestations.
+Our work addresses this gap by presenting the comprehensive, fine-grained dataset benchmark for multilingual, multicultural, and multievent online polarization, enabling robust cross-lingual and context-aware modeling.
 3 POLAR Dataset Construction
-We aim to construct a multilingual, multi-event, and multi-platform social media dataset spanning multiple continents.
-The dataset is designed to cover Indo-European and Afro-Asiatic language families, with a particular focus on balancing linguistic diversity and resource availability.
-It includes high-resource languages such as English, Spanish, and German; medium-resource languages like Arabic and Urdu; and low-resource languages such as Hausa and Amharic.
-While several Indo-European languages are represented (e.g., English, Spanish, German, Urdu), the dataset also incorporates key Afro-Asiatic languages (e.g., Hausa, Arabic, Amharic) to ensure broader geographical and linguistic representation.
-The following section will discuss the dataset construction in detail.
 3.1 Operational Definitions
-In this study, polarization refers to the process or phenomenon in which opinions, beliefs, or behaviors become more extreme or divided, leading to a greater distance or conflict between differing groups.
-We also classify the types of polarization as follows:
-  • Political/ideological polarization: This type of extremism focuses on division, intolerance, and conflict between political parties and followers.
-  • Racial or ethnic polarization: This type of polarization focuses on ethnic identity or racial origin and incites division, intolerance, and conflict between ethnic groups or races.
-  • Religious polarization: This type of polarization focuses on religious identity and incites division, intolerance, and conflict between religious followers
-  • Gender polarization: This type of polarization refers to the exclusion, discrimination, and marginalization of individuals based on their gender.
-  • Sexual orientation polarization: This refers to the increasing division and distinction between different sexual orientations within society, often leading to heightened tensions, misunderstandings, conflicts, or marginalization among various groups.
-  • Other: Polarized texts targeting other groups/identities such as economy class, technology, media, polarization, etc are classified as others.
-Furthermore, we define a polarized text manifestations or tactics if it exhibits one or more of the following characteristics:
+In this work, we define polarization as the increasing extremity of opinions, beliefs, or behaviors, resulting in heightened intergroup divisions and conflict. Polarization types are classified as:
+  • Political polarization: Focuses on division, intolerance, and conflict between political parties and followers.
+  • Racial or ethnic polarization: Focuses on ethnic identity or racial origin and incites division, intolerance, and conflict between ethnic groups or races.
+  • Religious polarization: Focuses on religious identity and incites division, intolerance, and conflict between religious followers
+  • Gender/ Sexual polarization: Refers to the exclusion, discrimination, and marginalization of individuals based on their gender or sexual orientations.
+  • Other: Polarized texts targeting other groups or identities not covered above, such as economic class, technology or media.
+In addition to topical categories, we further distinguish polarization by its rhetorical manifestations, defined as follows:
   • Stereotype: A generalized belief that attributes specific characteristics to all members of a group, often neglecting individual differences, thereby reducing complex personalities to simplistic and uniform representations.
@@
   • Dehumanization: The process of depriving a group or individual of their human qualities or personality by comparing them to animals, machines, or objects, or otherwise denying their humanity, dignity, or individuality.
-  • Extreme Language and Absolutism: The use of language that is extreme or makes definitive, all-encompassing statements, often involving words like "always," "never," "worst," or "best," and presenting issues in a dichotomous manner such as "us versus them" or "right versus wrong.
-  • Lack of Empathy or Understanding: The absence of compassion or recognition for other viewpoints or experiences in the text.
+  • Extreme Language and Absolutism: The use of language that is extreme or makes definitive, all-encompassing statements, often involving words like “always”, “never”, “worst”, or “best”, and presenting issues in a dichotomous manner such as “us versus them” or “right versus wrong”.
+  • Lack of Empathy: The absence of compassion or recognition for other viewpoints or experiences in the text.
   • Invalidation: The act of denying or dismissing the identity and existence of individuals or groups, thereby rejecting their sense of self and their presence.
-See Appendix A for further details and examples for each manifestation of polarization.
+Appendix E contains more details and examples for each manifestation of polarization.
+[^1] Since we define the above polarization manifestations as rhetorical tactics, we have used the terms “manifestation” and “rhetorical tactics” interchangeably.
 3.2 Data Collection
-We collected data from various online platforms, including X (formerly Twitter), Facebook, Reddit, Bluesky, Threads, and news/commentary forums.
-We use a dynamic keyword-driven strategy tailored for each language.
-Human experts curated keyword lists to reflect culturally and politically significant discourse across regions and events.
-Table˜1 illustrates the compositions, size, and event coverage of the dataset.
-The Amharic dataset focuses on the Northern Ethiopia or Tigray War in Ethiopia; Arabic texts cover a broad array of social and economic topics; English data centers on US elections and international conflicts; the German corpus features election discourse and migration debates; the Hausa subset captures religious and ideological discussions in West and Central Africa; Spanish content spans abortion, migration, and indigenous/gender rights; and Urdu texts reflect political and sectarian divides.
-Language Source(s) Event Train Dev Test Total Instances
-Amharic Facebook, X Tigray War in Northern Ethiopia 3500 500 1000 5000
-Arabic Facebook, X, Threads, News Social and economic issues 1482 212 424 2118
-English X, BlueSky, News US elections and international conflicts 2117 303 605 3025
-German X, BlueSky, Reddit Election discourse and migration debates 2426 347 694 3467
-Hausa Facebook, X Religious and ideological discussions in Africa 3893 557 1113 5563
-Spanish X, BlueSky Abortion, migration, and indigenous/gender rights 1400 201 401 2002
-Urdu X Political and sectarian conflicts 1960 280 560 2800
-Table 1: Summary of dataset composition: sources, sizes, and event coverage by task and language.
-3.3 Annotation Process
-Given the cultural and linguistic breadth of POLAR , we developed detailed, multilingual annotation guidelines and deployed a hybrid strategy combining crowdsourcing with trained community annotators.
-For crowdsourced annotators, we first filtered candidates based on their previous annotation experience and annotation agreement.
-Specifically, we selected those who had completed prior tasks with an agreement score of at least 0.8.
-We also conducted a pilot task consisting of 50 items to further filter annotators based on their agreement scores and average labeling time.
-For community annotators, we provided a training session followed by a pilot task to assess their performance before assigning the next batch of annotation tasks.
-During the annotation process, we continuously monitored annotator performance in a batch-wise manner to ensure that the overall process remained on track.
-Annotation Guidelines: We developed the guidelines in English and Amharic, and then translated and culturally adapted them for each target language.
-The English annotation guidelines are provided in Appendix A.
+Data Sources: We collected data from a range of online platforms, including major social media sites (e.g., X, Facebook, Reddit, Bluesky, Threads, YouTube comments, Weibo, and Zhihu) and local news or commentary forums (see Table 6 in the Appendix).
+For several languages (Chinese, Turkish, Polish, Burmese, and Italian), we sampled and re-annotated instances from existing toxic or hate speech datasets, including ToxiCN (Bai et al., 2025), COLD (Deng et al., 2022), the Turkish Hate Speech Dataset (Çöltekin, 2020), Myanmar Hate Speech (Kyaw et al., 2024), HaSpeede2 (Sanguinetti et al., 2020), HODI (Nozza et al., 2023), and BAN-PL (kołos2024banpl).
+Event Selection: We curated the dataset to cover diverse real-world events, grounding event selection in the sociopolitical and socioeconomic contexts specific to each language and cultural setting.
+The data span a broad range of events and issues, including armed conflicts (e.g., the Tigray War in Ethiopia, the Russia–Ukraine conflict, and the Gaza genocide), elections and party politics (e.g., the 2024 U.S.
+and 2025 German elections), public health crises, large-scale migration, climate change, and broader socioeconomic debates.
+The dataset also includes discussions related to gender and indigenous rights, religion, and ideology.
+For some languages, such as Bengali, a broader sampling strategy was adopted due to the lack of sufficiently large event-specific data on the selected platforms.
+# Lang. Inner Agr. ($\kappa$) Total PolarDetect PolarType PolarManifest
+Polarized (%) Political Racial / Ethnic Religious Polarization Gender / Sexual Other Stereo- type Vilifi- cation Dehuman- ization Extreme Language Lack of Empathy Invalid- ation
+1 eng 0.39 4,834 1,767 (37%) 1,726 422 168 108 190 730 1,272 586 1,156 536 879
+deu 0.10 4,771 2,274 (48%) 1,959 883 531 281 658 1,728 1,435 712 1,038 1,272 775
+urd 0.29 / 0.70 5,346 3,714 (69%) 3,603 2,908 2,954 2,739 2,713 3,328 3,460 2,973 3,324 3,007 3,059
+ben 0.59 5,000 2,127 (43%) 1,701 38 97 26 503 298 1,199 535 236 95 89
+hin 0.49 4,117 3,510 (85%) 3,051 500 2,417 472 540 2,047 2,683 750 2,082 2,336 2,703
+ori 0.46 3,552 1,021 (29%) 744 179 225 119 130 354 385 24 476 56 120
+nep 0.79 3,008 1,510 (50%) 518 422 239 158 354 806 947 198 816 318 450
+pan 0.55 2,609 1,280 (49%) 803 153 205 291 233 424 1,038 574 624 324 637
+ita 0.39 5,038 2,165 (44%) 412 926 368 461 219
+spa 0.26 4,958 2,479 (50%) 1,351 945 787 665 665 1,355 1,517 443 1,199 1,187 526
+rus 0.39 5,023 1,525 (30%) 696 494 205 284 119
+pol 0.46 3,587 1,504 (42%) 1,313 323 131 165 232
+fas 0.78 4,943 3,656 (74%) 2,170 120 476 296 1,197 649 2,850 213 835 487 394
+2 hau 0.48 5,477 587 (11%) 267 173 139 44 21 234 68 193 165 48 13
+arb 0.25 5,070 2,268 (45%) 1,205 874 424 553 847 1,691 1,896 555 1,540 863 411
+amh 0.59 4,999 3,747 (75%) 3,339 1,296 99 29 1,239 2,728 2,398 657 1,527 879 799
+3 zho 0.64 6,421 3,208 (50%) 376 1,475 127 1,085 552 1,931 1,188 323 522 506 307
+mya 0.13 4,334 2,508 (58%) 1,095 228 133 459 1,956
+4 swa 0.56 10,487 5,257 (50%) 279 3,721 371 234 833 4,160 4,324 1,340 2,509 3,120 2,456
+5 khm 0.83 9,960 9,042 (91%) 1,825 147 336 169 6,565 6,799 152 122 225 1,093 651
+6 tel 0.70 3,550 1,885 (53%) 766 603 318 471 842 398 781 88 477 933 809
+7 tur 0.46 3,566 1,776 (50%) 1,569 579 557 221 193 1,453 1,169 378 1,575 369 159
+Total 110,650 58,810 (53%) 30,768 17,409 11,307 9,330 20,801 31,113 28,762 10,664 20,326 17,429 15,237
+Table 1: Number of samples labeled as positive for each annotation task across languages.
+Inner agreement values denote inter-annotator agreement per language (Fleiss’s $\kappa$ unless otherwise noted).
+denotes exceptions: German uses Krippendorff’s $\alpha$; Punjabi reports identical Krippendorff’s $\alpha$ and Cohen’s $\kappa$; Urdu reports Fleiss’s $\kappa$ / Cohen’s $\kappa$.
+Polarization manifestation annotations are not available for Italian, Russian, Burmese, and Polish.
+Languages are ordered by language families and sub-branches as defined in Table 5.
+To collect data for these events, we adopted a dynamic, keyword-driven strategy tailored to each language and topic.
+Keyword lists were curated by human experts and native speakers to capture culturally and politically salient discourse and were used to retrieve data from online platforms.
+Table 7 in the Appendix provides an overview of the dataset composition, size, and event coverage.
+Dataset Quality Control: To ensure high-quality annotations across languages, we implemented several steps to ensure data quality throughout the quality control process.
+Before annotation, native speakers developed language-specific preprocessing pipelines.
+These pipelines included standard NLP procedures such as tokenization, word count filtering, and duplicate detection.
+Instances that were either too short or excessively long based on language-specific thresholds were removed.
+For anonymization, all usernames and URLs were replaced with standardized placeholders.
+For some languages, LLMs were used during the pre-filtering stage to increase the proportion of polarized content.
+3.3 Annotation Process and Guidelines
+We used a hybrid annotation strategy, leveraging crowd-sourced annotators and trained community annotators for low-resource languages where crowd-sourced annotation support is limited.
+For the crowd-sourced setting, annotators were selected based on their prior experience and annotation quality.
+Specifically, we filtered candidates using historical annotation agreement scores and conducted pilot rounds to identify those with consistent performance.
+Only annotators achieving a Fleiss’ Kappa score of at least 0.8 were retained for the main task.
+For community annotation, we recruited native speakers with at least a bachelor’s degree.
+Annotators received training, followed by a pilot round to assess their understanding and performance.
+Subsequent annotation was assigned in batches, with annotator performance monitored continuously to ensure consistency and adherence to guidelines.
+The trained annotators used POTATO Pei et al.
+(2022) and Label Studio [^2] https://labelstud.io as annotation platforms, while the crowd-sourced annotators used Mechanical Turk [^3] https://www.mturk.com and Prolific [^4] https://www.prolific.com.
+Annotation Guidelines: Given the cultural and linguistic breadth of POLAR, we developed detailed, multilingual annotation guidelines (see Appendix E) in English, and then translated and culturally adapted them for each target language.
 Annotators were instructed to:
-  • Identify whether a text is polarized;
-  • If the text is classified as polarized, tag the type of polarization (political, racial/ethnic, religious, gender/sexual identity, other);
-  • If the text is classified as polarized, tag its manifestations (stereotyping/deindividuation, vilification, dehumanization, extreme language, lack of empathy, invalidation).
+  • Identify whether a text is polarized
+  • If the text is classified as polarized, tag the type of polarization (political, racial/ethnic, religious, gender/sexual identity, other)
+  • If the text is classified as polarized, tag its manifestations or rhetorical tactics (stereotyping/deindividuation, vilification, dehumanization, extreme language, lack of empathy, invalidation).
 Multiple labels were allowed due to the conceptual and contextual overlap often observed in polarized content.
-Annotator Recruitment and Annotation Process: We used both expert and crowd-sourced annotation strategies.
-Trained annotators used POTATO and Label Studio, while crowdsourced annotators used Mechanical Turk and Prolific.
-To evaluate inter-annotator agreement, we report both Cohen’s Kappa and Fleiss’ Kappa.
-Since different teams were responsible for dataset creation across various languages, we provide a detailed description of the annotation process for each language:
-  • Amharic: We used trained annotators to annotated 5,000 samples in five batches using the POTATO annotation tool.
-Annotators completed three rounds of guideline training and received continuous feedback, with ongoing supervision to ensure quality control.
-The annotation quality achieved a Fleiss’ Kappa of 0.49, with full agreement in 69.62% of cases and a peak pairwise Cohen’s Kappa of 0.65.
-  • Arabic: Three topic-aware annotators labeled 2,118 samples using POTATO.
-Final polarization distribution was 16% positive, with Cohen’s Kappa at 0.296.
-Annotations were revised to ensure thematic consistency.
-  • German: Annotation was conducted through Prolific.
-Annotators were screened via surveys and underwent pilot testing.
-Final annotation used a locally hosted POTATO interface adapted for multilabel inputs.
-Inter-annotator agreement was moderate, with notable consistency on polarized cases.
-  • Hausa: We used trained annotators and Label Studio with iterative feedback and internal consistency checks. Data focused on religion and ideology on social media.
-  • Spanish and English: Amazon Mechanical Turk (MTurk) was used with strict quality filters, such as task completion history and approval ratings.
-Annotators received feedback and were contacted throughout for quality assurance.
-English annotation reached a Kappa of 0.52 (up from 0.31 after worker filtering); Spanish showed lower agreement with an average Kappa of 0.24.
-  • Urdu: Two approaches were used: (1) Manual Annotation: 1,792 samples labeled in batches with average Cohen’s Kappa of 0.55 (batch 1: 0.79; batch 2: 0.56; batch 3: 0.39).
-Disagreements were resolved by a fifth annotator.
-(2) Prolific: 1,460 samples were annotated in three batches.
-Initial agreement was low (Fleiss’ Kappa = 0.05).
-Additional filters (education, region) raised agreement in later batches to 0.28 and 0.30.
-The final dataset contains 2,800 high-quality Urdu samples.
-Figure 3: Percentage of polarization per language
-3.4 Dataset Characteristics
-Language Total Polarization Polarization Types Polarization Manifestations
-gender/sexual political religious racial/ethnic other vilification extreme_language stereotype invalidation lack_of_empathy dehumanization
-Amharic 5000 3753 29 3342 99 1297 189 2398 1527 2729 799 880 657
-Arabic 2118 338 102 178 86 171 220 413 341 283 240 175 102
-English 3025 1021 18 892 44 98 5 342 179 138 85 45 39
-German 3467 1596 241 1981 448 782 653 1371 1001 1420 1258 1124 568
-Hausa 5563 625 46 285 149 183 22 74 179 253 13 53 202
-Spanish 2002 1057 254 742 571 521 405 798 622 843 205 599 307
-Urdu 2800 1860 236 1533 557 423 126 1548 1409 979 670 632 501
-Table 2: Number of samples labeled positive for each annotation task across languages.
-Labels are grouped by task: Polarization (Task 1), Polarization Types (Task 2), and Polarization Manifestations (Task 3).
-We now analyze key trends and insights across the dataset based on annotated labels. Table 2 provides a detailed quantitative breakdown of positive labels for each annotation task.
-Polarization Prevalence: As shown in Figure 3 and further detailed in Table 2, Amharic (75%) and Urdu (66%) have the highest polarization rates, tied to their sources - Northern Ethiopia’s Tigray conflict and Pakistan’s political and sectarian climate.
-Spanish (53%), German (46%), and English (34%) fall in the mid-range.
-Arabic (12%) and Hausa (11%) reflect broader topical focus and possibly platform moderation effects.
-Hausa shows the lowest polarization rate despite covering sensitive religious and ideological topics, possibly due to moderation or cultural norms in expression.
-Datasets with acute sociopolitical events (e.g., Amharic, Urdu) show higher polarization.
-Topic scope, moderation, and platform dynamics significantly influence polarization prevalence.
-Types of Polarization: As illustrated in Figure LABEL:fig:types and Table 2, political polarization dominates in Amharic (67%), German (57%), Urdu (55%), and Spanish (37%).
-Racial/ethnic and religious types also appear prominently in Spanish, Amharic, and German.
-Gender/sexual identity polarization is generally low but visible in Spanish (13%) and Urdu (8%).
-Thematic emphasis (e.g., migration, elections) shapes type distribution.
-Spanish and Urdu show diverse forms of identity-based polarization, while political conflict dominates Amharic, Urdu, and German.
-Polarization Manifestations: Figure LABEL:fig:manifestation and Table 2 show the types of manifestations.
-Stereotyping/deindividuation and vilification are most frequent in Urdu, Amharic, German, and Spanish.
-Urdu’s vilification rate is highest (55%).
-Extreme language is strong in Urdu (50%), Spanish (31%), and Amharic (31%).
-Invalidation and dehumanization peak in German and Urdu.
-Conflict-heavy datasets (Urdu, Amharic) correlate with higher rates of hostile manifestations.
-Platform moderation and topic scope likely suppress such features in Arabic and Hausa.
-Together, these analyses highlight how linguistic, cultural, and contextual variables - alongside event salience and platform dynamics - shape the structure and tone of online polarization.
-The POLAR dataset offers a robust benchmark for examining these effects at scale.
+3.4 Annotators’ Reliability
+To evaluate annotation quality, we report Fleiss’ Kappa as the inter-annotator agreement (IAA) metric.
+As shown in Table˜1, the IAA scores vary between languages, with the majority showing moderate agreement and a few, such as "khm" and "tel" achieving good agreement.
+Although guidelines were standardized, their interpretation was influenced by cultural and political context, especially in languages with lower agreement, where some terms may not have direct equivalents across cultures.
+Latent content or sarcasm often required annotators to draw on their own socio-political knowledge, highlighting the perspectivist nature of polarization (Cabitza et al., 2023).
+Thus, low agreement can indicate socio-pragmatic complexity rather than error, signaling that polarization markers may not have universal meanings and that divergences can reveal inherent ambiguity in stimuli or interpretation (Aroyo and Welty, 2015).
+Examples illustrating such ambiguities are provided below.
+Unanimous polarized: “However I got tanned in Bodrum and I don’t wait in line at the hospital.” This sentence was rated as polarized by all three annotators with high confidence.
+Here, the sentence refers to perceived privileges granted to the immigrant community.
+Non-unanimous polarized: “Do you agree with the idea that new citizens in our country shouldn’t be allowed to vote for ten years?” This sentence did not receive consensus among annotators, leaning toward being polarized.
+The topic mentioned here is more complex and sensitive compared to the previous example.
+3.5 Dataset Statistics
+A comprehensive analysis of the dataset was conducted using the annotated labels to support systematic examination.
+Table 1 provides a quantitative breakdown of positive labels for each annotation task.
+For further details on data sources, language-wise composition, and polarization statistics, including distributions of polarized instances, types, and manifestations, see Appendix B (Tables 67 and Figure 4).
+Sample data is shown in Table LABEL:tab:sample_data, 10, and 11.
+# Lang. PolarDetect PolarType PolarManifest
+mBERT XLM-R RemBERT LaBSE twitter- roberta-hate afro-xlmr- large mBERT XLM-R RemBERT LaBSE twitter- roberta-hate afro-xlmr- large mBERT XLM-R RemBERT LaBSE twitter- roberta-hate afro-xlmr- large
+1 eng 74.83 76.34 77.51 77.27 79.59 75.54 31.29 27.33 43.17 34.68 36.99 19.09 41.83 44.50 47.90 46.58 49.04 47.82
+deu 65.69 69.03 69.81 68.57 65.97 59.14 49.34 48.89 54.23 53.26 49.39 51.34 43.74 45.24 47.46 47.13 44.28 45.18
+urd 71.60 68.74 75.62 74.20 63.45 73.51 68.76 72.16 72.96 73.91 71.05 67.16 72.13 72.83 74.75 77.09 72.49 43.00
+ben 79.03 83.68 82.74 82.50 36.81 83.04 24.61 16.59 10.14 26.82 10.14 13.62 21.08 24.12 5.05 25.83 9.59 5.64
+hin 76.37 75.06 45.98 78.23 61.80 65.49 61.40 63.58 73.30 69.97 38.83 57.87 53.71 69.12 72.69 70.13 63.58 62.07
+ori 41.72 41.72 71.14 77.60 42.42 68.73 19.44 28.14 32.92 42.27 19.23 29.47 8.37 4.84 2.78 24.00 5.77 9.47
+nep 85.04 84.04 87.48 88.26 69.86 76.09 45.75 48.78 70.53 66.42 32.69 49.16 55.15 53.92 58.11 60.19 39.19 35.55
+pan 69.98 66.50 73.57 74.91 61.03 62.28 29.59 28.22 40.88 38.73 27.32 29.08 42.07 40.07 45.80 48.35 38.57 43.31
+spa 70.36 71.27 75.24 74.99 71.96 74.24 58.22 54.54 65.01 59.34 57.78 58.64 41.17 43.34 50.63 48.04 43.76 46.45
+ita 54.35 34.50 60.93 56.40 53.00 47.78 21.48 21.83 25.99 26.13 21.08 23.47
+rus 70.20 71.61 77.39 74.77 42.33 74.80 31.97 21.93 47.33 40.94 15.81 36.44
+pol 74.45 74.54 77.96 77.26 68.45 67.62 31.63 31.29 46.84 41.55 24.65 24.48
+fas 80.09 78.34 77.50 83.50 54.64 80.89 46.69 47.21 54.18 52.46 27.72 27.75 35.53 34.42 41.35 39.51 24.72 34.78
+2 hau 82.97 47.19 75.91 82.09 78.32 81.82 21.44 15.75 15.53 21.58 16.68 2.76 19.97 16.43 21.44 20.55 18.29 22.16
+arb 78.24 78.43 81.12 81.23 68.37 82.20 49.49 50.43 58.89 55.89 33.40 56.16 49.94 53.05 55.53 56.76 43.21 56.30
+amh 42.45 68.83 72.60 76.43 49.46 58.28 29.32 39.48 24.27 47.09 28.19 28.43 41.21 26.94 42.47 51.16 41.28 45.75
+3 zho 83.96 86.45 87.08 86.44 72.49 85.83 57.39 47.83 69.35 63.12 34.05 45.17 40.50 37.91 39.09 46.10 27.72 38.87
+mya 84.10 82.44 84.33 86.07 70.35 82.31 42.62 34.54 25.55 55.06 26.05 42.52
+4 khm 47.58 47.58 76.41 73.70 47.58 68.38 23.06 50.12 65.79 58.60 15.89 17.20 14.54 29.77 35.60 34.28 14.57 14.26
+5 swa 76.54 76.02 78.17 79.04 75.93 78.77 36.39 33.20 43.35 40.18 35.63 38.46 55.69 54.32 56.21 56.51 53.79 54.82
+6 tel 84.86 79.63 88.93 88.93 65.16 78.62 41.16 32.36 43.68 42.56 30.77 36.73 37.87 21.55 39.79 39.24 29.65 35.83
+7 tur 69.80 74.47 70.94 75.02 65.42 73.06 36.55 43.92 51.03 48.41 30.76 31.33 36.92 44.83 44.48 44.92 36.83 44.36
+Table 2: Average macro-F1(%) scores for PolarDetect, PolarType, and PolarDetect  across languages and multilingual encoders.
+The best and second performance scores are highlighted in blue and orange respectively.
 4 Experimentation and Results
 4.1 Experimental Setup
-To evaluate POLAR , we conducted baseline experiments on three polarization detection tasks:
-  • Task 1: Binary classification - determining whether a text is polarized.
-  • Task 2: Multi-label classification - identifying polarization types (e.g., political, religious, ethnic).
-  • Task 3: Multi-label classification - detecting polarization manifestations (e.g., stereotyping, vilification, dehumanization).
-For data split, we used 70% for training, 10% for validation, and 20% for testing, as summarized in Table˜1. We pursued two main experimental paradigms:
-  1. 1.
-    Fine-tuning Multilingual Language Models (MLMs): We fine-tuned six multilingual models including InfoXLM (chi2021infosxlm), LaBSE (feng2022language), RemBERT (chung2020rethinkingembeddingcouplingpretrained), XLM-R (conneau2019unsupervised), mBERT (devlin2019bert), and mDeBERTa (he2021debertav3) for monolingual and crosslingual experiments.
-  2. 2.
-[…]
+To evaluate POLAR, we conducted baseline experiments on three polarization detection tasks: (1) classifying texts as polarized or not, (2) identifying polarization types, and (3) detecting polarization manifestations.
+For data splitting, we used 70% for training, 10% for validation, and 20% for testing, as summarized in Table˜7.
+All experiments were conducted using the EncouRAGe framework proposed by  Strich et al.
+(2025), which provides a standardized evaluation protocol for language model benchmarking.
+We benchmarked SLMs and LLMs.
+The list of evaluated models are listed below and Appendix C:
+  • Fine-tuning SLMs: We fine-tuned six SLMs, including four general-purpose models: mBERT (Devlin et al., 2019), XLM-R Conneau et al.
+(2020), RemBERT (Chung et al., 2021), and LaBSE (Feng et al., 2022).
+In addition, we evaluated two models build on social media and multilingual training corpus: twitter-roberta-hate (Antypas and Camacho-Collados, 2023), a RoBERTa-based encoder specialized for hate-speech detection, and AfroXLMR-large-76L (Adelani et al., 2023), an XLM-R variant optimised for African and other low-resource languages.
+  • Evaluating LLMs in Zero- and Few-shot Settings: We evaluated large language models in zero- and few-shot settings, including Qwen2.5-7B-Instruct, Qwen-3-8B, LLaMA-3.1-8B-Instruct, Ministral-3-14B-Instruct-2512, Gemma-3-27B-IT, GPT-4.1-Nano, and GPT-OSS-120B.
+For brevity, we refer to them as Qwen2.5, Qwen3, LLaMA3.1, Mistral3, Gemma3, GPT4.1 and GPT-OSS respectively.
+The exact models used are stated in the Appendix section C.
+The prompts used for LLM zero-shot and few-shot settings are shown in Appendix F.
+# Lang. PolarDetect PolarType PolarManifest
+Qwen2.5 Qwen3 LLaMA3.1 Mistral3 Gemma3 GPT4.1 GPT-OSS Qwen2.5 Qwen3 LLaMA3.1 Mistral3 Gemma3 GPT4.1 GPT-OSS Qwen2.5 Qwen3 LLaMA3.1 Mistral3 Gemma3 GPT4.1 GPT-OSS
+1 eng 71.19 69.42 69.17 70.67 79.84 77.19 77.43 72.45 66.76 69.08 73.79 74.37 69.59 77.73 63.33 62.38 60.62 63.91 64.18 63.39 68.33
+deu 60.02 58.76 61.53 61.34 73.86 70.89 69.72 66.61 61.26 64.17 67.31 71.71 63.63 70.04 57.74 60.10 56.82 58.71 57.93 59.41 65.18
+urd 40.97 48.46 58.55 51.26 75.88 72.02 70.42 42.03 44.24 49.27 45.11 50.65 46.92 49.32 42.20 50.32 52.14 54.74 53.22 48.44 52.46
+ben 65.62 59.75 68.90 69.83 80.85 76.15 80.28 68.96 66.10 65.71 71.05 74.04 72.64 77.13 58.47 54.54 51.23 50.47 51.89 56.99 60.34
+hin 43.94 36.53 48.52 45.67 67.13 59.04 58.96 65.32 54.48 59.96 68.80 75.12 69.73 71.68 48.09 53.62 55.74 60.50 58.43 53.04 58.53
+ori 53.35 50.14 49.30 42.38 65.03 58.39 67.20 65.87 64.74 65.11 48.61 75.81 66.68 71.29 55.64 55.22 54.27 49.76 59.46 57.88 61.66
+nep 49.03 53.48 65.01 61.84 85.28 74.16 82.68 69.31 67.29 65.09 73.78 80.95 72.30 82.29 62.67 67.08 61.88 70.01 73.34 69.59 76.30
+pan 40.69 40.42 49.80 41.64 71.95 71.67 65.92 56.51 54.20 56.98 54.21 68.20 61.73 66.19 51.10 54.90 53.20 55.22 56.73 58.48 61.94
+spa 66.64 68.25 70.42 70.04 71.43 73.64 76.91 67.69 66.07 61.82 69.05 69.81 65.93 72.18 60.00 59.86 57.46 57.30 59.84 60.34 66.36
+ita 66.24 63.20 67.52 67.36 73.72 72.08 74.60 68.48 64.68 63.59 70.19 73.27 68.47 74.70
+rus 73.77 71.01 71.12 72.88 70.57 72.85 74.36 70.26 66.33 57.78 73.43 69.88 68.47 77.94
+pol 57.15 58.33 63.46 57.25 68.88 65.42 76.94 70.56 62.91 65.26 68.87 75.88 70.42 79.46
+fas 32.17 28.44 40.10 29.71 54.79 48.71 51.61 59.53 52.96 63.23 54.93 71.17 60.74 66.33 58.62 55.68 56.52 55.40 56.63 58.98 62.93
+2 hau 51.40 49.25 55.37 50.44 46.36 52.18 55.02 55.41 53.65 50.69 54.59 54.04 56.84 59.03 51.01 47.28 43.09 43.56 43.66 48.43 48.33
+arb 65.22 58.82 70.02 65.47 80.70 77.72 78.87 63.22 62.77 62.62 66.17 71.22 64.39 72.25 62.74 64.57 62.64 67.04 67.97 65.02 73.93
+amh 25.16 26.97 39.14 23.50 74.47 50.58 57.00 52.85 51.55 61.67 47.26 77.01 61.20 67.31 50.78 53.99 56.51 53.43 59.77 55.98 61.52
+3 zho 58.20 59.70 70.57 62.62 82.48 78.51 80.52 77.38 73.32 68.54 75.32 80.63 74.76 81.15 69.99 67.72 58.16 64.35 65.82 70.44 77.90
+mya 40.63 48.31 51.68 37.72 78.40 58.41 59.96 52.65 55.76 58.48 49.23 67.78 58.09 59.48
+4 khm 10.56 8.94 11.04 8.65 13.32 11.25 12.53 50.51 46.30 57.32 45.32 55.21 48.71 49.43 47.96 50.04 50.14 49.28 49.30 48.66 48.50
+5 swa 40.50 46.76 57.42 52.29 65.64 62.00 66.15 61.31 57.52 61.63 62.87 69.49 66.34 68.81 52.94 54.04 54.63 55.11 56.89 56.22 60.41
+6 tel 42.99 44.23 43.02 36.82 42.97 45.23 43.51 52.81 53.88 55.82 52.24 57.88 52.70 50.95 50.78 54.96 54.08 55.68 55.29 56.59 52.13
+7 tur 64.11 58.86 66.68 66.23 76.69 74.54 78.97 67.43 63.41 62.19 70.38 75.41 70.27 77.49 60.38 61.54 60.84 62.94 65.08 62.73 71.20
+Table 3: F1-Macro resulting from the zero-shot LLM experiments with the POLAR dataset. The highest value per language is highlighted in blue.
 4.2 Results and Analysis
-Monolingual Cosslingual
-Task Lang. InfoXLM LaBSE RemBERT XLM-R mBERT mDeBERTa InfoXLM LaBSE RemBERT XLM-R mBERT mDeBERTa
-1. Polarization Amharic 77.96 78.92 81.93 80.64 53.71 74.98 23.26 49.44 5.63 0.00 0.00 83.21
-Arabic 58.39 61.94 67.19 70.42 52.48 54.68 29.68 35.87 27.76 2.78 37.79 25.63
-English 72.09 69.84 75.43 76.08 72.77 74.94 1.90 51.68 53.41 46.89 33.44 53.45
-German 29.75 63.46 67.50 67.78 58.45 64.38 0.62 51.97 64.59 41.87 46.05 25.94
-Hausa 59.57 66.41 67.43 66.41 59.32 65.62 8.21 22.92 6.47 1.50 6.25 25.13
-Spanish 38.69 64.04 70.98 57.00 59.31 54.16 45.70 68.66 67.26 21.09 68.34 62.17
-Urdu 1.07 66.35 79.95 43.91 65.63 60.68 1.60 30.36 0.00 0.00 3.68 68.67
-2. Types Amharic 24.22 38.13 43.65 25.71 17.93 25.00 11.56 20.65 2.49 0.00 0.00 26.76
-Arabic 22.97 40.23 42.12 37.58 27.53 35.22 11.49 23.73 8.78 2.11 12.47 7.93
-English 16.04 23.25 31.38 24.10 21.07 17.70 10.84 19.69 16.42 12.22 11.70 3.93
-German 13.52 58.58 61.19 58.56 54.13 40.64 12.03 35.86 29.59 9.72 23.98 11.88
-Hausa 18.56 19.14 17.38 18.09 19.61 18.87 3.20 9.97 3.91 1.39 5.78 5.90
-Spanish 43.26 66.07 67.76 58.07 57.94 43.40 7.89 47.06 15.38 0.43 32.02 14.60
-Urdu 27.14 51.94 51.60 45.38 38.02 33.13 3.77 13.62 6.33 0.00 3.94 20.30
-3. Manifestations Amharic 43.52 47.56 47.63 43.17 33.18 43.29 15.57 27.07 8.64 0.00 0.00 43.58
-Arabic 40.05 51.55 52.52 55.61 42.18 47.73 16.56 30.68 22.14 0.00 19.57 17.24
-English 14.40 15.01 19.39 18.61 18.60 15.15 7.16 10.16 10.05 5.62 10.69 8.85
-German 38.49 49.88 52.74 51.70 46.85 51.91 2.38 36.12 27.05 0.00 23.38 12.93
-Hausa 19.23 20.04 19.18 18.89 18.74 18.93 5.74 5.86 3.77 3.12 6.19 6.43
-Spanish 38.94 50.00 51.04 45.02 45.17 35.09 2.34 40.63 11.83 0.40 35.99 23.56
-Urdu 34.26 52.20 53.64 41.32 45.90 47.01 2.10 19.16 11.54 0.00 1.88 48.58
-Table 3: Average F1-Macro for all three tasks.
-In the monolingual settings, we train and evaluate the model for each language separately.
-In the crosslingual setting, we train on all languages within a language family (AfroAsian: Amharic, Arabic, Hausa, and Urdu), Germanic: English, German, and Spanish) except the target language, and evaluate on the test set of the target language.
-The best performance scores are highlighted in blue and orange, respectively.
-4.2.1 Monolingual Setup Results
-To benchmark the performance of state-of-the-art multilingual language models (MLMs) for polarization detection, we fine-tuned six prominent pretrained models: InfoXLM (chi2021infosxlm), LaBSE (feng2022language), RemBERT (chung2020rethinkingembeddingcouplingpretrained), XLM-R (conneau2019unsupervised), mBERT (devlin2019bert), and mDeBERTa (he2021debertav3).
-Each model was evaluated in a monolingual (within-language) setting across three tasks: (i) binary polarization classification, (ii) multi-label polarization types classification, and (iii) multi-label polarization manifestations detection.
-For Task 1, binary polarization classification, RemBERT consistently achieved the highest macro F1-scores for most languages, while XLM-R performed best in Arabic, English, and German.
-In the multi-label polarization types classification (Task 2), RemBERT again provided the top scores for Spanish, except for Hausa and Urdu, where mBERT and LaBSE performed best, respectively.
-For Task 3, multi-label polarization manifestations classification, the results were generally lower across all models, with RemBERT delivering the highest scores for most languages, except for Arabic and Hausa.
-4.2.2 Cross-Lingual MLM Results
-For cross-lingual transfer experiments, each MLM was trained on all other languages within the same language family (Afro-Asian or Germanic), excluding the target language, and evaluated on the target language’s test set for the same three tasks.
-In the cross-lingual setup for polarization classification (Task 1), mDeBERTa achieved the highest F1-macro for Amharic, English, Hausa, and Urdu.
-For Arabic, German, and Spanish, mBERT, RemBERT, and LaBSE performed slightly better than the other models, respectively.
-For Task 2, polarization types classification, performance was generally lower as expected in the cross-lingual scenario.
-LaBSE yielded the best results for four languages, while mDeBERTa performed better for Amharic and Urdu.
-Similarly, for Task 3, polarization manifestation classification, the results were quite low overall, but LaBSE and mDeBERTa showed relatively better performance compared to the other models.
-This suggests that general polarization detection transfers more easily across languages than the more nuanced identification of polarization types and manifestations, which are culturally and linguistically specific.
-Effective cross-lingual transfer likely requires aligned annotation schemes and fine-tuning strategies sensitive to these differences.
-4.2.3 Zero and Few-shot LLM Performance
-We evaluated seven state-of-the-art LLMs (Qwen3-8B, LLaMA-3.1-8B, Mixtral-8x7B, GPT-4o, GPT-4o-mini, GPT-4.1-mini, and Llama-4-Scout-17B-16E-Instruct) across languages.
-The prompts used for evaluation are provided in the Appendix B.
-We explicitly provided definitions of tasks and leveraged the models’ chain-of-thoughts capabilities by encouraging reasoning before labeling.
-In the few-shot setting, we provided 4 example instances for each task.
-To assess the impact of prompt language, we also compared model performance when prompted in English versus the respective target language.
-As Table  4 illustrates, the closed-source models GPT-4o, GPT-4o mini, and GPT-4.1 mini, consistently outperform the other four LLMs, demonstrating superior performance in polarization recognition.
-However, their performance on Hausa remains notably low, highlighting both the challenge of understanding low-resource languages and the value of our multilingual dataset in exposing the limitations.
-In contrast, smaller open-source models, such as Qwen3-8B, LLaMA-3.1-8B, and Mixtral-8x7B, achieve lower F1-Macro scores and perform inconsistently across languages.
-Their F1-Macro scores improve by 5 10% when provided with few-shot examples.
-However, the larger LLaMA model, LLaMA-4-scout-17B, demonstrates performance comparable to GPT-4o in polarization classification.
-An interesting phenomenon is observed: Smaller models show improved classification performance in the few-shot setting, while larger models such as the GPT series and LLaMA-4-17B do not consistently gain additional improvement with few-shot examples.
-This suggests that few-shot learning offers advantages to smaller models than to larger ones.
-Task Lang GPT 4o GPT 4o-mini GPT 4.1-mini Mistral-7B Qwen3-8B Llama-3.1-8B Llama4-17B
-Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot Zero-shot Few-shot
-1. Polarization Amharic 71.30 68.74 55.40 58.04 63.52 63.20 44.84 48.24 55.60 65.51 50.98 58.84 69.96 67.91
-Arabic 71.51 79.65 58.21 68.41 76.20 79.58 33.65 53.06 65.95 70.76 61.51 67.21 73.58 77.45
-English 75.46 79.56 64.78 80.30 77.43 77.48 53.78 70.00 69.09 66.86 66.55 69.90 76.90 75.48
-German 72.04 71.12 63.03 68.49 73.77 69.60 56.17 61.77 65.42 68.76 59.90 65.40 70.89 70.35
-Hausa 34.73 44.87 28.21 34.07 39.45 48.06 19.82 28.15 45.91 49.18 38.73 40.80 47.85 54.53
-Spanish 72.69 65.76 62.51 67.96 71.28 69.22 51.24 61.00 65.42 68.76 60.20 62.53 66.11 66.27
-Urdu 72.25 78.83 59.19 69.92 73.15 72.63 48.27 54.79 66.99 70.89 55.61 59.77 73.93 68.50
-2. Types Amharic 39.67 36.59 33.8 34.89 32.49 34.44 44.21 63.24 63.07 75.90 78.26
-Arabic 35.00 39.30 37.18 36.82 33.07 36.77 49.39 53.91 53.22 52.86 55.03
-English 34.96 35.67 33.79 38.73 35.22 38.22 48.28 68.50 67.37 62.71 67.09
-German 58.86 58.58 54.36 57.94 52.70 55.15 44.10 71.82 70.62 73.89 76.20
-Hausa 18.96 19.11 16.90 18.15 17.00 18.86 43.34 67.24 70.47 71.75 74.66
-Spanish 48.98 52.93 52.82 55.09 48.89 49.76 43.15 65.96 67.87 69.82 71.43
-Urdu 55.6 59.64 53.97 54.78 51.82 55.09 48.40 63.36 64.83 60.01 61.52
-3. Manifestations Amharic 44.09 45.59 40.01 44.78 40.53 42.49 0 0 0 0 0
-Arabic 51.90 56.90 46.96 51.01 52.57 60.76 0 0 0 0 0
-English 22.95 25.87 20.43 25.78 28.41 26.67 0 0 0 0 0
-German 42.84 45.16 43.33 53.78 37.62 41.23 0 0 0 0 0
-Hausa 8.58 8.14 8.42 8.37 8.55 9.33 0 0 0 0 0
-Spanish 41.00 40.99 39.85 46.33 38.82 40.66 0 0 0 0 0
-Urdu 52.89 54.43 48.25 58.17 48.34 51.89 0 0 0 0 0
-Table 4: F1-Macro resulting from the zero- and few-shot LLM experiments with the POLAR  dataset. The highest value per language is highlighted in blue.
+SLMs Models: Table 2 presents the results of six small language models.
+Overall, RemBERT and LaBSE show comparable performance across most languages, achieving the best or the second best macro-F1 scores.
+RemBERT is designed to balance representation across high-, mid-, and low-resource languages, while LaBSE relies on bilingual sentence-level alignment between English and low-resource languages.
+These training strategies enhance their ability to mid- and low-resource language understanding, which is reflected in their consistent improvements over mBERT and XLM-R, particularly for languages such as Amharic, Odia, Italian.
+The twitter-roberta-hate model is finetuned on English Twitter dataset covering emoji, stance, hate speech, and emotion.
+This training boosts its performance on English polarization detection, polarization types classification and polarization manifestation recognition.
+AfroXLMR-large-76L is finetuned upon XLM-R on African languages datasets.
+As a result, it performs well on African language in our dataset, including Arabic, Hausa, and Swahili.
+Polarization detection is comparatively easier for multilingual BERT-based models, which achieve relatively high macro-F1 scores.
+In contrast, recognizing polarization types (politics, gender, racial, religious, and others) is substantially more challenging.
+Performance drops even further for polarization manifestation recognition (stereotype, vilification, dehumanization, extreme language, lack of empathy, invalidation), where macro-F1 scores decrease markedly.
+This gap highlights the limitations of current models in capturing fine-grained and implicit polarization manifestations and the importance of POLAR  for advancing research on nuanced polarization understanding.
+LLMs Performance: The Tables 2, 3, and 4 present an overall picture of model performance in polarization detection, types, and manifestations across the languages, using the Macro-F1 metric as a measure of accuracy.
+Across all experiment setups, models consistently achieve their highest scores in Polarization Detection, followed by the classification of Polarization Types, while identifying Polarization Manifestations remains the most challenging task for both encoder models and LLMs likely due to the latent nature of how polarization is expressed, requiring semantic understanding that state-of-the-art models find difficult to generalize across diverse cultural contexts.
+Table 2 presents the the performance of encoders (mBERT, XLM-R, RemBERT, LaBSE, twitter-roberta-hate, and afro-xlmr-large), demonstrating significant stability across languages, especially in the PolarDetect task.
+On the other hand, Table 3 presents a more mixed picture through zero-shot LLM experiments.
+While high-resource languages like English (eng) and Chinese (zho) maintain high detection scores (reaching 79.84% and 82.48% respectively), the models exhibit a significant drop in performance for languages with unique scripts or less resources.
+For example, Khmer (khm) detection scores drop to a range between 8.65% and 13.32%, indicating that without prior examples, these generative models may not generalize.
+Table 4 demonstrates the transformative impact of few-shot prompting, where providing the LLMs with a few examples significantly improves performance for many languages.
+For instance, Urdu (urd) has the detection score rise from 40.97% in the zero-shot setting to 73.81% with GPT4.1 in the few-shot setting.
+In this few-shot setting, LLMs also outperform BERT-family encoders in the complex PolarManifest task, for example, with Arabic (arb) reaching a peak of 75.49%.
+In summary, while BERT-family models remain the most efficient for binary detection, LLMs with few-shot prompting show potential for handling more complex, fine-grained classification tasks when sufficient examples are provided.
+Finally, some of the linguistic families exhibit clustered behaviors.
+For instance, a language cluster emerges among South Asian languages like Bengali (ben), Hindi (hin), Nepali (nep), and Oriya (ori), as these languages exhibit similar performance improvements in the few-shot LLM experiments, compared to zero-shot settings (Table 4).
+Nevertheless, the most dominant predictor of model performance seems to be resource-tier of the language, which often aligns with geographic and economic regions.
+# Lang. PolarDetect PolarType PolarManifest
+Qwen2.5 Qwen3 LLaMA3.1 Mistral3 Gemma3 GPT4.1 GPT-OSS Qwen2.5 Qwen3 LLaMA3.1 Mistral3 Gemma3 GPT4.1 GPT-OSS Qwen2.5 Qwen3 LLaMA3.1 Mistral3 Gemma3 GPT4.1 GPT-OSS
+1 eng 77.98 77.79 74.90 75.82 79.82 76.81 79.74 72.17 73.52 70.48 77.33 76.96 74.01 79.24 62.37 63.40 61.56 63.41 64.35 63.77 69.40
+deu 65.49 65.79 65.93 64.34 71.26 71.66 69.61 65.03 67.75 68.37 69.58 72.44 68.31 69.61 58.36 60.76 55.97 60.59 53.72 57.59 65.03
+urd 57.89 55.42 55.48 43.96 62.62 73.81 66.26 40.80 49.04 47.02 45.40 51.85 48.14 49.39 43.73 57.34 54.55 59.44 58.42 49.75 56.67
+ben 66.84 63.00 59.91 65.61 79.64 78.89 78.32 70.14 72.51 71.14 73.37 75.11 71.12 76.71 57.01 53.29 51.41 53.12 53.72 58.90 60.36
+hin 52.92 43.45 41.42 52.13 58.54 62.61 56.26 59.85 54.86 55.54 69.68 71.31 66.72 70.99 47.77 51.70 54.76 59.41 58.65 54.29 57.49
+ori 60.85 57.23 60.81 43.60 63.62 65.71 66.09 63.98 66.26 67.26 50.38 75.68 67.56 73.38 54.71 56.56 53.24 48.69 59.73 57.57 62.57
+nep 64.38 74.27 64.25 69.50 84.64 82.44 81.20 68.49 73.98 71.72 77.16 81.57 74.45 82.38 59.81 64.29 60.89 68.35 70.57 72.03 76.92
+pan 51.19 60.47 56.38 52.95 69.77 72.07 69.76 61.04 64.38 62.86 64.38 69.43 66.76 70.57 50.94 53.07 53.95 55.51 57.22 58.66 62.82
+spa 68.01 70.43 67.35 66.91 72.15 74.05 74.61 68.46 68.96 64.16 70.36 71.67 68.65 73.39 59.49 58.69 58.10 59.80 60.00 62.54 66.29
+ita 59.37 55.83 48.57 51.98 65.92 71.82 72.02 60.67 63.55 63.15 64.49 70.46 68.51 72.98
+rus 74.96 73.70 62.74 70.59 74.32 74.04 77.45 69.78 70.74 63.57 74.10 73.97 71.19 78.71
+pol 58.07 62.86 53.14 59.10 68.38 71.54 79.17 70.56 66.92 69.71 71.19 76.91 72.50 79.79
+fas 40.77 41.32 40.45 39.32 54.58 54.22 50.53 57.14 59.51 66.76 59.26 69.72 62.67 65.54 58.08 55.41 54.25 56.35 57.53 59.24 62.47
+2 hau 55.01 53.92 55.60 50.47 56.24 55.08 59.28 57.89 55.73 55.49 54.61 57.79 56.69 60.64 50.80 49.01 47.69 46.19 45.98 50.87 49.94
+arb 75.54 72.88 73.42 72.21 79.58 78.37 78.40 64.36 65.53 66.16 70.07 71.44 70.50 72.89 64.89 66.55 61.96 67.07 67.6 67.58 75.49
+amh 38.51 35.70 37.66 28.59 74.58 51.81 54.67 50.41 59.07 64.44 55.74 76.65 59.64 66.07 50.90 55.80 51.50 53.36 58.72 52.91 61.37
+3 zho 56.69 68.17 56.79 66.71 69.81 78.22 74.57 71.13 78.00 72.03 76.32 76.87 79.12 77.00 69.97 67.39 61.25 62.78 63.98 72.01 76.47
+mya 62.17 71.12 60.51 50.96 73.90 68.46 63.07 56.23 63.35 68.11 58.24 71.63 65.54 63.89
+4 khm 24.70 13.66 15.75 12.98 18.90 16.52 15.69 53.99 51.21 62.92 51.81 57.24 51.99 51.41 49.45 50.55 49.55 52.19 50.47 48.27 50.11
+5 swa 53.67 58.43 58.75 59.20 67.70 64.56 66.61 61.14 65.17 65.74 66.71 70.80 67.59 69.81 56.27 54.88 53.13 53.58 56.04 57.62 61.13
+6 tel 58.52 56.59 53.40 54.56 60.02 59.74 52.15 52.15 55.13 56.16 56.64 57.24 56.24 53.92 53.79 55.79 55.41 54.61 59.03 56.71 52.53
+7 tur 72.29 68.12 66.23 68.70 79.62 76.45 78.28 64.02 66.43 62.74 71.08 76.03 69.09 77.67 60.74 58.18 60.60 64.70 64.83 64.14 72.87
+Table 4: F1-Macro scores(in %) from the few-shot LLM experiments with the POLAR  dataset. The highest value per language is highlighted in blue.
 5 Discussion
-Our multilingual and cross-cultural investigation into online polarization reveals that polarization is a deeply contextual and event-driven phenomenon.
-Its prevalence and intensity are closely tied to the sociopolitical climate, public discourse, and digital environment of each region.
-In societies experiencing major crises, such as the war in Ethiopia, religious tensions in Pakistan, and international conflicts like the wars in Ukraine and Gaza that influence the US election, we find polarization to be especially amplified on online platforms.
-Conversely, broader topical selection, moderation, or more diffuse societal tensions are associated with noticeably lower polarization.
-The types of polarization evident in online discourse are shaped by the most salient local issues: political (e.g., in Ethiopia, the US, and Germany), religious (e.g., in Pakistan), ethnic (e.g., in Ethiopia), or gender-based (e.g., in Spain) divides tend to dominate where these subjects are at the forefront of public debate or conflict.
-This underscores the importance of accounting for each region’s unique sociopolitical history and cultural landscape when analyzing or intervening in digital polarization.
-Polarization also manifests in varied rhetorical forms, including stereotyping/deindividuation, vilification, and exclusionary or hostile language.
-These patterns are particularly pronounced amid contentious issues and persistently divided communities, such as civil war, election, migration issues, religious divide, and so on.
-From a computational perspective, our results demonstrate both the feasibility and the challenges of detecting polarization in multilingual settings.
-Binary polarization classification is relatively tractable for well-resourced languages and clear-cut contexts, especially when leveraging state-of-the-art multilingual models (Sec.
-4).
-However, classifying the specific types and manifestations of polarization introduces significantly greater ambiguity and requires a much deeper, context-sensitive understanding.
-Performance on these more nuanced tasks remains limited.
-This highlights important directions for future work, including the integration of cultural signals and context or event understanding, the development of more robust multilingual embeddings, and the pursuit of consistent annotation (with a great deal of annotator training and follow up, specially for the crowdsourcing annotations).
-The wide variability across languages, contexts, and events further affirms the necessity of designing language, culture, and event specific approaches, particularly for underrepresented regions in global NLP research.
-Taken together, these findings highlight that effective detection and mitigation of online polarization must move beyond generic or monocultural solutions.
-There is a pressing need for context-aware, adaptable methodologies that acknowledge both the universal and the local characteristics of polarized discourse, ensuring relevance and effectiveness across the world’s diverse digital landscapes.
+Performance within the same language family:
+It is not necessary that models uniformly achieve a high performance across all languages within the same language family.
+For example, while a model may perform strongly on Chinese, its performance on Burmese still lags behind, demonstrating that high-resource language performance does not guarantee transferability to related lower-resource languages such as Burmese.
+Few-shot vs. Zero-shot:
+Few-shot performance is not always superior to zero-shot results.
+For larger models, such as Gemma3 and GPT-OSS, the few-shot setting underperforms zero-shot, indicating that large LLMs are inherently capable of recognizing polarized sentences without requiring in-context exemplars.
+In contrast, for smaller models, including Qwen2.5, Qwen3, LLaMA3.1, and Mistral3, in-context examples (3-shot prompting) can improve performance, showing that smaller LLMs benefit more from in-context learning.
+LLMs vs. SLMs:
+SLMs demonstrate improved performance in detecting polarization after fine-tuning, but still struggle to distinguish polarization types and manifestations, reflecting limited semantic knowledge of domain-specific concepts.
+Conversely, LLMs generally exhibit a stronger understanding of social-science polarization constructs, enabling better recognition of polarization types and manifestations, yet they remain weaker in direct polarization classification tasks.
+This suggests that LLMs possess richer implicit knowledge of social-science constructs, whereas SLMs currently lack comparable semantic grounding for fine-grained polarization distinctions.
+6 Error Analysis: Misclassification Cases
+Based on our error analysis (See detail in Appendix G Table 8), we identified a misalignment between the model’s classification logic and human judgment of polarization.
+The model relies on a deterministic, surface-level heuristic: it classifies text as polarized only when it detects explicitly named and opposed groups within the sentence (e.g., "Israeli forces vs.
+Palestinians").
+Conversely, if the text expresses hostility but names only one group or relies on implied opposition, the model defaults to labeling it as non-polarized.This error stems from models’ reliance on textual pattern alone, while human annotators draw upon cultural and contextual knowledge to interpret hostility and implicit group conflict.
+7 Implications
+Theoretical Implications: Our dataset highlights the complexity of online polarization, emphasizing its deep cultural and contextual nature.
+It reveals the current limits of NLP models in detecting implicit rhetorical tactics, underscoring the need for culturally-aware frameworks.
+Practical Implications: The dataset provides a valuable benchmark for developing and evaluating models capable of detecting nuanced forms of online polarization across multiple languages and contexts.
+It supports the creation of more culturally sensitive and robust tools for the monitoring and mitigation efforts of online discourse.
+Methodological Implications: Our multi-label, multi-platform annotation approach underscores the importance of culturally sensitive and detailed labeling strategies.
+The variability of model performance across languages and contexts indicates a pressing need for methods that integrate cultural signals, multimodal data, and contextual embeddings to improve robustness and reduce performance disparities in social NLP applications.
 8 Conclusion
-In this study, we introduced POLAR, the first multilingual, multicultural, and multievent dataset for benchmarking online polarization.
-Our findings showed that polarization was deeply context-dependent, manifesting in diverse types and rhetorical forms shaped by local sociopolitical dynamics.
-Theoretically, we provide new evidence that polarization is a multifaceted, cross-lingual phenomenon strongly influenced by language, culture, and current events, rather than being uniform across settings.
-On a practical level, our benchmark revealed that while binary detection of polarization was feasible with current models, accurately identifying specific types and manifestations remained a persistent challenge, especially for multilingual and low-resource contexts.
-This work established a foundation for further computational studies of digital polarization and for developing more culturally robust moderation and intervention tools.
-Looking ahead, future research should integrate richer contextual and cultural signals into model architectures, refine annotation guidelines for cross-regional consistency, and expand evaluations to include additional languages, social domains, and further events.
-By releasing our data and benchmarks, we aimed to catalyze further innovation toward nuanced detection and effective mitigation of online polarization worldwide.
+In this study, we introduced POLAR, a comprehensive, multilingual, and multi-event dataset designed to advance the understanding and detection of online polarization across diverse linguistic and cultural contexts.
+By annotating over 110,000 instances along three critical dimensions, we created a nuanced, fine-grained resource.
+This dataset captures the complex rhetorical tactics and social dimensions that underpin polarized discourse.
+Our extensive benchmarking of SOTA SLMs and LLMs reveals that while current models are reasonably effective at binary polarization detection, they face significant challenges in accurately identifying polarization types and rhetorical manifestations, especially in low-resource and culturally nuanced settings.
+These findings emphasize the deep contextual and implicit nature of online polarization and highlight the limitations of existing NLP approaches.
+Importantly, our work underscores the critical need for culturally aware, adaptable, and context-sensitive models to effectively monitor and mitigate digital polarization globally.
+The resources and benchmarks provided herein aim to catalyze future research, fostering the development of more inclusive and robust tools for analyzing social phenomena in multilingual and multicultural online environments.
 Limitations
@@
 Despite rigorous protocols, labeling polarization remains subjective; we encourage responsible, ethically grounded use of this resource and discourage misuse.
-As Table 5 demonstrated, the inter-annotation agreement scores range from 0.2 to 0.6, corresponding to fair to moderate levels of agreement according to established benchmarks (landis1977kappa).
-This level of agreement further highlights the inherent difficulty of polarization recognition and underscores the necessity and value of our proposed work.
-Language Annotation Platform Kappa Score
-Amharic Potato 0.65
-Arabic Potato 0.296
-English MTurk 0.52
-German Prolific 0.29
-Hausa Potato 0.48
-Spanish MTurk 0.24
-Urdu Potato 0.55
-Table 5: Dataset annotation platforms and inter-annotator agreement scores
-Appendix A Annotation Guidelines
+Appendix A Language and its language family
+The POLAR dataset covers 22 languages from seven linguistic families. The coverage language and their corresponding branches is presented below:
+# Lang. ISO-639 Language Family Sub-branch
+1 English eng Indo-European Germanic
+German deu Indo-European Germanic
+Urdu urd Indo-European Indo-Aryan
+Bengali ben Indo-European Indo-Aryan
+Hindi hin Indo-European Indo-Aryan
+Odia ori Indo-European Indo-Aryan
+Nepali nep Indo-European Indo-Aryan
+Punjabi pan Indo-European Indo-Aryan
+Spanish spa Indo-European Romance
+Italian ita Indo-European Romance
+Russian rus Indo-European Slavic
+Polish pol Indo-European Slavic
+Persian fas Indo-European Iranian
+2 Hausa hau Afro-Asiatic Chadic
+Arabic arb Afro-Asiatic Semitic
+Amharic amh Afro-Asiatic Semitic
+3 Chinese zho Sino-Tibetan Sinitic
+Burmese mya Sino-Tibetan Tibeto-Burman
+4 Khmer khm Austroasiatic Mon-Khmer
+5 Swahili swa Niger–Congo Bantu
+6 Telugu tel Dravidian Dravidian
+7 Turkish tur Turkic Turkic
+Table 5: Language covered and its language families
+Appendix B Data Statistics
+Table 6 summarizes the distribution of instances across data sources in the POLAR dataset.
+The data primarily originate from major social media platforms, with Twitter contributing over half of the instances.
+Additional data are drawn from news websites, online forums, and other social platforms to ensure coverage of diverse discourse contexts.
+A smaller portion of the data comes from existing datasets that were re-annotated to align with our polarization framework.
+Source No. of Instances %
+X (Twitter) 58,984 53.3%
+News/websites 15,208 13.7%
+Bluesky 8,099 7.3%
+YouTube 7,159 6.5%
+Reddit 5,705 5.2%
+Existing Dataset 4,794 4.3%
+Weibo 4,550 4.1%
+Facebook 2,928 2.6%
+Zhihu 1,088 1.0%
+Threads 852 0.8%
+Tieba 783 0.7%
+Wikipedia 500 0.5%
+Total 110,650 100%
+Table 6: Data sources of the POLAR dataset.
+B.1 Dataset Composition by Language
+Table 7 presents a language-wise overview of the POLAR dataset, including data sources, targeted events or topics, and inter-annotator agreement.
+For each language, data collection was tailored to platform availability and sociopolitical relevance, resulting in variation in event focus and discourse type.
+Inter-annotator agreement is reported primarily using Fleiss’s kappa, with alternative reliability measures noted where applicable.
+B.2 Polarization Statistics
+Figure 3 shows the distribution of polarization types, and Figure 4 illustrates the distribution of polarization manifestations across languages.
+Figure 3: PolarType  by Languages, for each language, the numeric value shows the count of instances assigned to a given polarization type, while the percentage reflects its share of the total annotated instances for that language.
+Figure 4: PolarManifest-by Languages.
+For each language, the numeric value shows the count of instances assigned to a given polarization manifestation, while the percentage reflects its share of the total annotated instances for that language.
+Language Data Source(s) Events/Topics Focused
+Amharic (amh) Facebook, X (Twitter) The Tigray War
+Arabic (arb)
+Existing datasets, Facebook,
+News, Reddit, Threads,X (Twitter)
+Social issues regading politics and religion
+Bengali (ben) YouTube comments Social discourse, generic contemporary topics
+Burmese (mya) Existing datasets Kyaw et al. (2024), Wikipedia
+Social issues regading politics, ethnicity
+and popular culture
+Chinese (zho) Tieba, Weibo, Zhihu
+Social issues regarding racism,
+sexuality/gender and religious discrimination
+English (eng) Bluesky, Local news, X (Twitter) US elections and international conflicts
+German (deu) Bluesky, Reddit, X (Twitter) COVID-19, and contemporary social issues
+Hausa (hau) Facebook, X (Twitter)
+Social issues regading politics, ethnicity
+and religion
+Hindi (hin) Bluesky, Reddit, X (Twitter)
+Social issues regarding politics, religion,
+and caste
+Italian (ita) YouTube, X (Twitter)
+Pride parade, immigration crisis
+, crime news, Italian justice reform
+Khmer (khm)
+Facebook, Specialised websites,
+Wikipedia, YouTube, Local news
+COVID-19, and contemporary social issues
+Nepali (nep) Facebook, Local news, X (Twitter) Social discourse, generic contemporary topics
+Odia (ori) Bluesky, Local news, X (Twitter) Social discourse, generic contemporary topics
+Persian (fas) Bluesky, X (Twitter) Social discourse, generic contemporary topics
+Polish (pol) Bluesky, Existing dataset kołos2024banpl COVID-19, and contemporary social issues
+Punjabi (pan)
+Existing dataset (pnbTenTen)
+https://www.sketchengine.eu/pnbtenten-western-punjabi-corpus ,
+Youtube Comments,
+X (Twitter)
+Social issues regarding politics, religion,
+and caste
+Russian (rus)
+Bluesky, X (Twitter),
+COVID-19 chatter dataset Banda et al. (2021)
+Social discourse, generic contemporary topics
+Spanish (spa) Bluesky, X (Twitter)
+2010’s inmigration movement,
+"Salvemos las dos vidas" movement,
+social issues regarding politics
+and gender inequality
+Swahili (swa) X (Twitter) Kenyan elections
+Telugu (tel) Facebook, Reddit, X (Twitter) Social discourse, generic contemporary topics
+Turkish (tur) X (Twitter), Existing dataset Çöltekin (2020) Social discourse, generic contemporary topics
+Urdu (urd) X (Twitter) Social discourse, generic contemporary topics
+Table 7: Summary of dataset composition regarding data sources and events or topics focused.
+Appendix C SLMs and LLMs Used
+C.1 Multilingual Encoders
+C.2 LLMs
+Appendix D Experiment Settings
+D.1 Experiments settings
+For SLMs, we performed language-specific fine-tuning for 3 epochs using a learning rate of 2e-4.
+For LLMs in the few-shot setting , we use three shots.
+ALL prompts are written in English, while the in-context examples are provided in the target language.
+The full prompts is reported in Appendix F.
+All experiments were conducted on local GPU and implemented with EncouRAGe (Strich et al., 2025) framework to execute experiments and measure results.
+Appendix E Annotation Guidelines
 Annotation Guidelines
@@
   4. 2.4
-    Gender Polarization – Discrimination based on gender
+    Gender/Sexual Polarization – Discrimination based on gender or sexual orientation
   5. 2.5
-    Sexual Orientation Polarization – Tension or marginalization based on sexual orientation
-  6. 2.6
     Other – e.g., based on economic class, technology, media, etc.
-Note: Multiple types may be selected if applicable.
+Note: Select all relevant options if the text contains more than one of the above.
 Task 3: Manifestations of Polarization (if “Yes” in Task 1)
@@
   4. 3.4
-    Deindividuation – Ignoring individual identities Example: “All migrants are thieves.”
+    Extreme Language or Absolutism – Using words like “always”, “never”, or “worst”. Example: “We can never trust them.”
   5. 3.5
-    Extreme Language or Absolutism – Using words like “always,” “never,” or “worst” Example: “We can never trust them.”
+    Lack of Empathy or Understanding – Dismissing others’ perspectives Example: “Wearing the hijab is extremist.”
   6. 3.6
-    Lack of Empathy or Understanding – Dismissing others’ perspectives Example: “Wearing the hijab is extremist.”
-  7. 3.7
     Invalidation – Denying a group’s identity or existence Example: “There is no nation called Palestine.”
-Thank you for your valuable contribution to this research.
-Appendix B Prompt for Text Classification
-Task 1: Polarization Detection – Zero Shot
-Given the following texts in [X] language, classify them as polarized or non-polarized based on the following definition:
-Definition of Polarization: Polarization refers to the phenomenon where opinions, beliefs, or behaviors become more extreme or divided, resulting in increased separation and conflict between different groups.
-These groups may be based on politics or ideology, ethnicity or race, religion, gender, or sexual orientation.
-Polarization involves negative attitudes such as stereotyping, denigration, dehumanization, de-individualization, or intolerance toward individuals or groups outside one’s own, while exhibiting blind support and solidarity toward those within the group.
-Reasoning: Provide the reason why the text demonstrates polarization.
-Label: Return True if the text is polarized, otherwise False. Do not return any other value.
-Task 1: Polarization Detection – Few Shot
-Same instructions as above. Examples provided:
-  • Text: Black immigrants should leave the USA.
-    Reasoning: This statement stereotypes a group based on race and immigration status, implying a negative attitude and exclusion.
-    Label: True
-  • Text: Africans are lazy because they are backward.
-    Reasoning: Negative trait assignment based on ethnicity, showing polarization.
-    Label: True
-  • Text: Muslims are terrorists.
-    Reasoning: Promotes harmful stereotypes and dehumanization.
-    Label: True
-  • Text: African women are beautiful.
-    Reasoning: Positive generalization, not polarizing.
-    Label: False
-Task 2: Polarized Text Typing – Zero Shot
-Classify each polarized text into one or more of the following types:
-  • Gender/Sexual: Discrimination based on gender or sexual orientation
-  • Political: Division or conflict between political ideologies
-  • Religious: Conflict between religious groups
-  • Racial/Ethnic: Conflict or division based on race or ethnicity
-  • Other: Related to economy, technology, media, etc.
-Output: A list of five binary values in the order: [Gender/Sexual, Political, Religious, Racial/Ethnic, Other]
-Few Shot Examples
-  • Text: Black immigrants should leave the USA.
-    Output: [0, 0, 0, 1, 0]
-  • Text: Muslims are terrorists.
-    Output: [0, 0, 1, 0, 0]
-  • Text: Women should be confined to traditional roles.
-    Output: [1, 0, 0, 0, 0]
-  • Text: The economy is failing because of government policies.
-    Output: [0, 1, 0, 0, 0]
-Task 3: Polarization Manifestations – Zero Shot
-Analyze and identify the following manifestations in each text. Output a list of six binary values in this order:
-[Vilification, Extreme Language, Stereotype, Invalidation, Lack of Empathy, Dehumanization]
-Definitions
-  • Stereotype/Deindividuation: Generalizes traits, ignores individuality.
-  • Vilification: Defames a group or person, inciting fear.
-  • Dehumanization: Strips humanity, compares people to animals or objects.
-  • Extreme Language: Uses absolutes like always, never, us vs. them.
-  • Lack of Empathy: Dismisses others’ experiences or views.
-  • Invalidation: Denies the identity or existence of a group.
-Few Shot Examples
-  • Text: Black immigrants should leave the USA.
-    Output: [1, 1, 1, 1, 1, 0]
-  • Text: Africans are lazy because they are backward.
-    Output: [1, 1, 1, 1, 1, 0]
-  • Text: Muslims are terrorists.
-    Output: [1, 1, 1, 1, 1, 1]
-  • Text: African women are beautiful.
-    Output: [0, 0, 0, 0, 0, 0]
-Appendix C Dataset Samples
+Note: Select all relevant options if the text contains more than one of the above.
+Appendix F Prompts for Text Classification
+F.1 Prompts for polarization detection
+TASK: Determine whether the following {{lang_full}} text is polarized or not polarized.
+DEFINITION: Polarization refers to the process or phenomenon in which opinions, beliefs, or behaviors become more extreme or divided, leading to a greater distance or conflict between differing groups.
+Attitude polarization is the negative attitude that individuals or groups display towards individuals and groups outside their group while also showing blind support and solidarity towards people within their group.
+Polarization denotes stereotyping, vilification, dehumanization, deindividuation, or intolerance of other people’s views, beliefs, and identities.
+Speeches and articles that are shared on social media that incite division, groupism, hatred, conflict, and intolerance are classified as containing polarization.
+Only texts that clearly reflect attitude polarization should be classified as such.
+If a text includes one or more of the specified characteristics, it is classified as polarized.
+Conversely, social media texts that do not display any of these characteristics are classified as non-polarized.
+Always consider the context and the overall meaning of the text, not just individual words or phrases.
+INSTRUCTIONS: - Analyze the text carefully.
+- Classify the polarization of the given text.
+Use the given definition of polarization.
+- Provide a short reason explaining your decision.
+The reason should be English.
+- Assign a binary label: If the text is polarized, polarization is set to 1.
+If the text is not polarized, polarization is set to 0.
+- Do not include any text or formatting outside the JSON.
+OUTPUT FORMAT:
+{"reason": "the reason of the classification decision", "polarization": 0 or 1}
+Examples (if applicable):
+{% if examples %}
+{{examples}}
+{% endif %}
+Input Text:
+{{user_prompt}}
+F.2 Prompt for polarization type
+TASK: Determine whether the following {{lang_full}} text is polarized or not polarized and which type of polarization.
+DEFINITION: Polarization refers to the process or phenomenon in which opinions, beliefs, or behaviors become more extreme or divided, leading to a greater distance or conflict between differing groups.
+Polarization denotes stereotyping, vilification, dehumanization, deindividuation, or intolerance of other people’s views, beliefs, and identities.
+Only texts that clearly reflect attitude polarization should be classified as such.
+If a text includes one or more of the specified characteristics, it is classified as polarized.
+Conversely, social media texts that do not display any of these characteristics are classified as non-polarized.
+Always consider the context and the overall meaning of the text, not just individual words or phrases.
+TYPES OF POLARIZATION: - Political/ideological polarization: Political polarization refers to political beliefs and affiliations becoming more extreme.
+- Racial or ethnic polarization: This type of polarization focuses on ethnic identity or racial origin and incites division, intolerance, and conflict between ethnic groups or races.
+- Religious polarization: This type of polarization focuses on religious identity and incites division, intolerance, and conflict between religious followers.
+- Gender/Sexual polarization: This type of polarization refers to the exclusion, discrimination, and marginalization of individuals based on their gender and sexual orientations within society, often leading to heightened tensions, misunderstandings, or conflicts.
+- Other: polarization texts targeting other groups/identities such as economy, technology, media, polarization, etc.
+INSTRUCTIONS: - Analyze the text carefully, and classify the polarization of the given text and if the text is polarized the types of polarization.
+- Use the given definition of polarization and their types.
+Provide a short reason explaining your decision.
+The reason should be English.
+- Assign a binary label for polarization: If the text is polarized, polarization is set to 1.
+If the text is not polarized, polarization is set to 0.
+- Classify the text for the different types of polarization.
+- The answers need to be in the following order: [political/ideological, racial/ethnic, religious, gender/sexual, other].
+- The text can contain more than one type of polarization.
+If the text is not polarized set all polarization types to 0.
+Do not include any text or formatting outside the JSON.
+OUTPUT FORMAT:
+{   "reason": "the reason of the classification decision",
+    "polarization": 0 or 1,
+    "polarization Types": [0/1, 0/1, 0/1, 0/1, 0/1]
+{% if examples %}
+**EXAMPLES**:
+{{examples}}
+{% endif %}
+**TEXT**:
+{{user_prompt}}
+F.3 Prompt for polarization manifestation
+TASK: Determine the manifestation type(s) of polarization expressed in the following {{lang_full}} text.
+LABEL SET: ["stereotype", "vilification", "dehumanization", "extreme_language", "lack_of_empathy", "invalidation"] DEFINITIONS: - stereotype: Generalizes traits to all members of a group, ignoring individual differences.
+- vilification: Defames or demonizes a group or person through exaggerated or biased framing.
+- dehumanization: Denies humanity by comparing people to animals, machines, or objects.
+- extreme_language: Uses absolute or hyperbolic language (e.g., “always”, “never”, “us vs.
+them”).
+- lack_of_empathy: Shows indifference or refusal to understand others’ perspectives.
+- invalidation: Denies or rejects the identity, legitimacy, or existence of a group or individual.
+INSTRUCTIONS: - Select all labels from the list that apply (one or more).
+- If the text is not polarized, set labels = [0].
+- Provide a short reasoning sentence explaining your decision.
+- Do not include any text or formatting outside the JSON.
+OUTPUT FORMAT:
+    "reason": "the reason of the classification decision",
+    "polarization": 0 or 1,
+    "polarization Types": [0/1, 0/1, 0/1, 0/1, 0/1, 0/1]
+{% if examples %}
+**EXAMPLES**:
+{{examples}}
+{% end if %}
+**TEXT**:
+{{user_prompt}}
+Appendix G Error examples
+Text Gold Type Reason given by model
+The Cybersecurity Infrastructure Security Agency has frozen its election security work is reviewing everything it has done to help state local officials secure elections for the past 8 years.
+…The move accommodates Ts false claim of election fraud.
+@URL 0 FP The text criticizes a political figure’s claim as false and portrays a government agency as accommodating that claim, expressing a negative attitude toward the out-group and contributing to partisan division."
+Four key facts Hezbollah is not an Islamic sect or denomination.
+It is a terrorist org founded in 1985.
+Hezbollah is an Iranian Shia Islamist proxy based in Lebanon.
+Not a Lebanese resistance movement.
+0 FN The text vilifies Hezbollah by labeling it a terrorist organization and dismissing its claimed legitimacy, showing hostile stereotyping toward the group, which meets the criteria for attitude polarization.
+this new optic the President use that term lightly posted on X of him as a Kingis he serious?
+Are we no longer a Democracy?
+We have an election wearly voting starting tomorrow.
+Im running.
+Did King Donny stop elections?
+I appreciate your courage.
+1 FN The text merely questions the President’s rhetoric and expresses concern about democratic processes without demeaning, vilifying, or stereotyping any group, so it does not show attitude polarization.
+Black Officers Awarded \$16M After White Colleague’s Racial Comments 1 FN The headline reports an incident involving race but does not contain demeaning language, vilification, or calls for hostility toward any group, so it is not polarized.
+And thats not ethnic cleansing how ?
+1 FN The brief question does not contain explicit negative or vilifying language toward any group, nor does it promote division or intolerance, so it lacks clear attitude polarization.
+Table 8: Example texts, the human annotated gold labels, the error type (FN=False Negative, FP=False Positive) and the Reason given by the model
+Appendix H Dataset Samples
+The dataset samples are available at the following link: https://anonymous.4open.science/r/trial-data-05C6/README.md.
 Table 9: Examples of polarized statements with annotated types and manifestations.
 Types: P: Political, Re: Religious, Ra:Racial; Manifestations: V: Vilification, EL: Extreme Language, S: Stereotype, I: Invalidation, LE: Lack of Empathy, D: Dehumanization.
-Text T2 T3
+Text Types Manifestation
 "Denazification" as an excuse to explain Russia’s deployment of troops to Ukraine, alleging that Hitler was of Jewish ancestry P V, EL
@@
 A very moral army. IDF Gaza Israel P V, LE
-Text Lang Label
+Text Lang Polarized
 Gibt sogar recht viele Ex-Muslime. Alles sehr vernünftige Leute. DE TRUE
@@
 Table 10: German samples with labels and English translations
-Text Lang Label
+Text Lang Polarized
 y vos un random judio que comenta pelotudeces para poder figurar en algun lado SPA TRUE